A Complete Bibliography of ACM Transactions on Interactive Intelligent Systems (TIIS)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

25 June 2024 Version 1.34

Title word cross-reference

2 [GSG^{+}24]. 3 [HJ11, TSSL15, TRO^{+}19]. N [MLD^{+}21].

-generAItor [SKS^{+}24].

2017 [CDF19]. 2018 [BBQ20]. 2019 [BCKC20].

AAR [DKI^{+}21, KDA^{+}22]. AAR/AI [DKI^{+}21, KDA^{+}22]. Abilities [CLRT20, SCC14]. Ability [HKR^{+}18, MHC^{+}18]. Abnormal [MWP^{+}21]. Abstract [SYS^{+}15]. Academic [DMGG13]. Accelerating [RRS^{+}22]. Acceptability [KMO22]. Access [CSM^{+}12]. Accessible [MGP^{+}22]. Accountable [TVA^{+}21]. Accuracy [HKR^{+}18, KB17, SFKL20]. Accurate [PCNB17, SLVV23]. ACM [BBQ20, BCKC20, CDF19, KGCC18]. Acquisition [FBS13]. Action [DKI^{+}21, LPBG19]. Actionable [SML^{+}23]. Actions [ZFM^{+}24]. Activation [SVK16]. Active [BHS^{+}21, HJ11, HRAR18, KPSK17, SD23]. Activities [NKP^{+}19, SD23]. Activity [APY^{+}20, BBT^{+}15, DNV^{+}23, RNA^{+}23, YSD15, YK16, YVA^{+}24]. Ad [SMF^{+}23]. ADAM [HWM^{+}22]. Adaptation [SKS^{+}24]. Adapting [SVK16]. Adaptive [BLIC21, CMTB15, GSC^{+}17, HTOB^{+}20, HWM^{+}22, KDMA12, Moh21, PSOC16, SLH^{+}21, TRO^{+}19, VBCC20, WZB^{+}16, WES16, YDL^{+}17, ZLG22, HBP^{+}12, LPBG19]. Added [RAMZ^{+}17]. Addressing [WDN24]. adjusting [KKS22]. Adult [BPUP22, KMO22]. Adults [RSvO^{+}22, RD21, SBN^{+}18]. Adversarial [LWFM23]. Advertisements [PW19]. Advice [LDLW23, SHY16, TOF16]. Advising [DMGG13, HWM^{+}22]. Advisor [SHY16]. Aerial [DM17]. Aerobic [MVH20]. Aesthetic [BSKC21]. Affect [BDO^{+}16, CLP^{+}14, KD20, YLVE12, EWS12]. Affect-Aware [KD20]. Affectionate [CNI15]. Affective [MDR^{+}22, NCWC17, SYS^{+}15, CRP^{+}12, ONTHS11, DG12]. After [DKI^{+}21]. After-Action [DKI^{+}21]. Agent [BMD^{+}15, COAR16, DHS^{+}17, EBB^{+}22, MDR^{+}22, RSvO^{+}22, TSN^{+}16, YLC23, JIKB22, KCM22]. Agents [CRO^{+}16, HZKW21, JIKB22, KW16, KLC^{+}22, MGP^{+}22, MJJT24, OPM17, PLBW17, PAGM15, SBN^{+}18, VPB16, ZMLY19, BSBC12, INN13]. Aggregate [SFKL20]. Agile [CSO^{+}22]. agnostic [MNL^{+}23]. AI [DKI^{+}21, CSO^{+}22, CHST23, CWZF24, DKI^{+}21, JJT24, KDA^{+}22, LDM23, LDLW23, MJJT24, MVH20, MZR21, NSA^{+}22, NRB^{+}22, PBF^{+}23, PDA^{+}21, RFR^{+}24, SF23, Shn20, VPAJ^{+}23, VOA^{+}22, WY22, WWB^{+}23, Yan21, ZFM^{+}24]. AI-Assisted [WY22, RFR^{+}24]. AI-Powered [CWZF24]. AI-Supported [VOA^{+}22]. AID [SF23]. Aided [DM21]. Air [AUMT16, GGC^{+}19, HCD^{+}20, SLVV23]. AI’s [KDA^{+}22]. Algorithm [LBC^{+}21]. Algorithmic [DDGMP20, SW20]. Allow [RAMZ^{+}17]. aloud [FLT20]. Alphabet [PLBG19]. ALVA [KPSK17]. am [EBB^{+}22]. Amateur [SYS^{+}15]. Ambiguity [CDK^{+}18]. Ambiguous [WDN24]. among [MDR^{+}22]. Analogous [KCTH19]. Analysis [BMD^{+}15, BEH22, EVB^{+}20, FWY^{+}18, GSG^{+}24, KB17, LZL^{+}23, LC18, MBR16, MWP^{+}21, NKP^{+}19, PSC^{+}16, SYS^{+}15, SDW^{+}18, SKKM15, vdGPOvB^{+}13]. Analytic [CQW^{+}14]. Analytics [AGH^{+}23, DSEV19, DPS^{+}19, HMO22, JLH^{+}23, KPSK17, LWFM23, SMM20, SKG^{+}18, TVA^{+}21, WDN24, ZTSH23]. Analyzing [KKL^{+}19]. AnchorViz [SGR^{+}20]. Android [BSKC21]. animation [GCDN11]. Annotated [MKW21]. Annotation [BMD^{+}15, KP18, RFR^{+}24]. Announcement [ZB23]. Anomaly [Riv14]. Answering [SV22]. Anxiety [EBB^{+}22]. anxious [EBB^{+}22]. Anything [SG12]. AOI [LBC^{+}21]. AOI-shapes [LBC^{+}21]. application [KPP14]. Applications [EVB^{+}20, KTD17, NRB^{+}22, PCNB17, SML^{+}23, ZBGD14]. Approach [BFY^{+}21, COAR16, DSV17, DPS^{+}19, FB18, HBKM20, NSMM18, OPM17, PW19, SJS^{+}21, SMM20, SD23, TTL^{+}22, EWS12, ONTHS11]. Approaches [AGH^{+}23, BHS^{+}21, SGS^{+}12]. Appropriate [MJJT24]. Architecture [RPRK^{+}21, VLP^{+}15]. Archives [HMO22]. Areas [LBC^{+}21]. Argumentation [DMGG13]. Argumentative [RK16]. Arguments [RK16]. Armband [PLBG19]. Arousal [MKW21]. Art [GRNW15, SYS^{+}15, TFT21]. Article [Zho22]. articles [CSM^{+}12]. Artificial [GGK^{+}21, JJT24, TOF16, Zho21, BSBC12]. Artistic [DDGMP20]. Arts [SHA^{+}15b, SHA^{+}15a, YDL^{+}17]. Aspects [DDGMP20]. Assessing [COAR16]. Assessment [RSPT^{+}23]. Assistance [GRC^{+}12, HBP^{+}12]. Assistant [BPUP22, HWM^{+}22, MBD^{+}15]. Assisted [WY22, RFR^{+}24]. Assisting [YNK^{+}22]. Associated [DLMC15]. assuring [ZTD^{+}24]. Asynchronous [HC16]. attachment [HCDRB12]. attachment-like [HCDRB12]. Attacks [LWFM23]. Attention [DPA20, NWP^{+}18, PAGM15]. Attentive [BLIC21]. AttentiveVideo [PW19]. Attitude [CRO^{+}16]. Attributes [AL23]. Audience [YDL^{+}17]. Audio [EGQ16]. audiovisual [SGS^{+}12]. Auditory [FB18]. augmentation [MGF13]. Authentication [AUMT16]. authoring [YIS^{+}14]. Autism [WZB^{+}16]. Auto [FJZ^{+}22]. Auto-Icon [FJZ^{+}22]. Automated [BMD^{+}15, FJZ^{+}22, SF23, SHY16, ZTSH23]. Automatic [DGG16, DK15, FLT20, MBR16, PSOC16, SGS^{+}12]. Automatically [BDO^{+}16]. Automating [AMN^{+}13]. Automation [CLS^{+}18]. Autonomous [BEC^{+}20, HWM^{+}22]. AutoTutor [DG12]. Avatars [HHC20]. Aware [BFY^{+}21, BMD^{+}15, KK13, KD20]. Awareness [SF23, INN13].

back [DG12]. Backgrounds [NRB^{+}22]. Bad [ZFM^{+}24]. Bandit [BHNP20]. Bandit-Based [BHNP20]. Based [BHNP20, GGC^{+}19, NSMM18, OPM17, PASR^{+}16, SSL^{+}19, SG12, TSSL15, TVCL16, WHLZ18, AF22, AL22, BSKC21, CYW19, MJJT24, RAE^{+}14, SCLH21, SD23, WRH18, YLVE12, ZTD^{+}24]. Bases [KH12]. Basis [NYYT16, RK16, SLH^{+}21]. Bayes [KSW^{+}11]. Bayesian [MDC^{+}24]. Be [IOKY16a, YYI^{+}16]. Behavior [BLV14, CRO^{+}16, CNI15, DK15, IOKY16a, KMAM20, Moh21, RSvO^{+}22, RK16, SHA^{+}15b, SHA^{+}15a, SWS21, YLC23, YT15, CRC^{+}12, HCDRB12, JFNY13]. Behavioral [CMPS17, RHS19]. Behaviors [BLIC21, LC18, RAMZ^{+}17, YIS^{+}14]. Being [DPA20, MDR^{+}22]. Benefits [TSSL15]. Best [ZB23]. Between [Shn20]. Beyond [KB17, ZGZ^{+}16]. Beyond-Accuracy [KB17]. Bi [KKL^{+}19]. Bi-Level [KKL^{+}19]. Bias [FBS13, GSC^{+}17, PKCS12, RDP^{+}24]. Biased [WWB^{+}23]. Big [CMPS17, CGHS17]. Binary [NLN^{+}14]. Biohazard [PASR^{+}16]. Blocks [ZF17]. Body [NKP^{+}19, PSOC16, WES16, BSBC12, SDD12]. BONNIE [SB21]. Bottom [SCLH21]. Bottom-up [SCLH21]. Boxes [TVCL16]. Brain [YDL^{+}17].

Brain-Adaptive [YDL^{+}17]. Bridging [Shn20]. Brief [SHA^{+}15b]. browser [TDBL13]. Building [SB21].

Can [ZFM^{+}24, KK12]. Capabilities [KW16]. Capable [CNI15]. Captions [ACMH23]. Capturing [GRC^{+}12, KLR^{+}18]. Care [YYI^{+}16]. caregiving [HCDRB12]. Case [DSEV19, KLC^{+}21, SG12, VOA^{+}22, WWB^{+}23]. Case-Based [SG12]. casual [JFNY13]. Category [Zho22]. Center [DPA20]. Centered [FG18, FB18, Zho21, BHS^{+}21, PBF^{+}23, Shn20, Yan21]. Centred [SV22]. Change [Moh21]. Changes [DLMC15, MKS^{+}17, YLVE12]. Characterization [OBG^{+}19]. Characterizing [WBL^{+}13]. Chatbots [KMO22, SP22]. Child [RHS19]. children [ONTHS11]. Choices [EDR24]. Chronodes [PCK^{+}18]. Citation [WSL^{+}18]. Citizen [HCD^{+}20]. Classification [AHJR18, DGG16, DLMC15, JWK14, KPSK17, ZCC21, KSW^{+}11]. Classroom [BSKB21]. Classrooms [BDO^{+}16]. Client [MWP^{+}21]. Clinical [PBF^{+}23]. ClioQuery [HMO22]. Clouds [KD20]. Clustered [ZCC^{+}18]. Clustering [NSMM18, SMM20, WDN24]. CNN [DPA20]. Co [DHS^{+}17, JLH^{+}23, PBF^{+}23]. Co-Creative [DHS^{+}17]. Co-design [PBF^{+}23]. Co-Occurrences [JLH^{+}23]. Coach [MVH20]. Coaching [Moh21, ZT22]. Code [CWZF24, FJZ^{+}22, LZL^{+}23]. Coding [CDK^{+}18]. Cognition [Moh21]. Cognitive [BEC^{+}20, CLRT20, NCWC17, RPRK^{+}21, SCC14, ZLG22, CRC^{+}12]. cognitively [DG12]. Collaboration [DHS^{+}17]. Collaborative [FBC14, LCKM23, PYDM16]. Collection [ACMH23]. Collective [NWP^{+}18, PCK^{+}13]. College [WWB^{+}23]. Colonoscopy [VOA^{+}22]. Column [Zho20, Zho21]. Combating [GSC^{+}17]. Combinations [WDN24]. Combining [CLRT20, RSPT^{+}23]. Commerce [BHNP20]. Commodity [ZGZ^{+}16]. Common [DJH^{+}12, GRC^{+}12, KH12, LH12, Moh21]. Communicating [MHC^{+}18]. Communication [KCTH19, TSN^{+}16, SFTI^{+}13]. Communities [CWZF24]. Community [HCD^{+}20, VSR12]. Companion [CLP^{+}14]. Companionable [SBN^{+}18]. Comparing [CLRT20]. Comparison [PLBG19, TRO^{+}19, SGS^{+}12]. Comparisons [MSD^{+}16, WY22]. Competent [RSvO^{+}22]. Completion [CLJ21]. Completions [BLO22]. Complex [GGK^{+}21, TVCL16]. Complexity [GGK^{+}21]. Complying [SHY16]. Components [BHNP20]. Composition [LDS^{+}22]. Comprehensive [HMO22]. Computational [LZL^{+}23, OBG^{+}19, PKCS12]. Computer [And12, BDO^{+}16, ITB19, KCTH19, KKS22, PBKS20, SDD12]. Computer-Enabled [BDO^{+}16]. Computer-Mediated [KCTH19]. computers [DG12]. Computing [Yan21]. Concept [SGR^{+}20, RSPT^{+}23]. Conceptual [BEH22, GZGW23, WRH18]. Confidence [SLMK18]. Conflict [Hol18]. Connecting [GRNW15]. Consciousness [RSPT^{+}23]. Consideration [DLMC15]. Considering [BEH22]. Constraint [LZW^{+}24]. Constraints [BLO22]. Contact [XZY16]. Content [DSEV19, KB16, RAE^{+}14]. Content-based [RAE^{+}14]. Context [AL20, BMD^{+}15, CLP^{+}14, DPA20, HK16, HTOB^{+}20, KK13, KLR^{+}18, SVK16, SMF^{+}23, TSSL15, YT15]. Context-Aware [BMD^{+}15, KK13]. Context-dependent [HTOB^{+}20]. Context-Sensitive [CLP^{+}14, SMF^{+}23]. Contextual [YVA^{+}24]. Contextualization [GSC^{+}17]. Contingent [WZB^{+}16]. Continuous [SDD12, VOA^{+}22, EWS12, ZBGD14].

Control [GKP15, KK12]. Controllable [DBTV20]. Conversation [CSD14, AC13]. Conversational [EBB^{+}22, HZKW21, KLC^{+}22, MDR^{+}22, OPM17, PAGM15, SCKOGP15, SMF^{+}23, SP22, INN13, JFNY13]. Conversations [HBZ23, HC16, OBG^{+}19, ZF17, SGS^{+}12]. Conveys [YYI^{+}16]. ConVisIT [HC16]. Convolutional [LWFM23]. Cooperative [MDC^{+}24, MGF13]. Core [SMF^{+}23]. Core-Queries [SMF^{+}23]. correlation [GGC^{+}19]. Coverage [KB17]. Create [OPM17]. Creating [KK12, SBN^{+}18]. creation [PCK^{+}13]. Creative [DHS^{+}17]. Criminal [HZKW21]. Critical [AYL^{+}22]. cross [SFTI^{+}13]. cross-lingual [SFTI^{+}13]. Crowd [LC18, ZCC^{+}18]. Crowds [GRNW15]. Crowdsourced [MTML20]. Crowdsourcing [ACMH23, DAW18, SFKL20]. Cube [CCM14]. Cues [Hol18, JJT24]. Culture [CLS^{+}18]. Curious [RPRK^{+}21]. Current [BPUK18]. Customer [OBG^{+}19]. Customizable [HBP^{+}12]. Cyberbullying [DJH^{+}12]. Cyclicity [DM17].

D [HJ11, TSSL15, TRO^{+}19]. Daily [DFR15, SD23]. Dance [ABJ15]. Data [ACMH23, BHS^{+}21, CMPS17, CGHS17, CLRT20, FLJA23, GGC^{+}19, HBKM20, HMH^{+}19, JLH^{+}23, KTD17, LZL^{+}23, PBKS20, SDW^{+}18, SKKM15, SLH^{+}21, SCC14, SGR^{+}20, TTL^{+}22, WSL^{+}18, ZCC21, GCDN11, SGS^{+}12, GSG^{+}24]. Data-Driven [HBKM20, PBKS20, GCDN11]. Data-Record [FLJA23]. Dataset [MKW21]. Datasets [AGH^{+}23, HK16]. Debiased [WWB^{+}23]. debugging [KSW^{+}11]. decade [JR12]. Decision [CdGF^{+}13, DMGG13, DM21, GGK^{+}21, HSS^{+}21, PBF^{+}23, SLMK18, WY22]. Decision-making [HSS^{+}21]. Decisions [CHST23, KK13, TVA^{+}21, HBP^{+}12]. Decomposition [GSG^{+}24]. Deep [KCC^{+}19, RNA^{+}23]. Defects [BSKC21]. defined [BLO22, LBC^{+}21]. Degree [HBZ23]. Delivery [SF23]. dependent [HTOB^{+}20]. Depicting [SWS21]. Depression [EBB^{+}22]. Design [AUMT16, DK18, FB18, HC16, JJL16, KCM22, MDC^{+}24, MZR21, VBCC20, YLC23, YIS^{+}14, ONTHS11, PBF^{+}23, TDBL13, ZBGD14]. Designing [HBKM20, LDM23, MVH20, Moh21, MNL^{+}23]. Designs [FJZ^{+}22]. Detect [BDO^{+}16]. Detecting [NCWC17]. Detection [BSKC21, DPR^{+}22, DK15, DJH^{+}12, FLT20, KP18, MR19, MWP^{+}21, Riv14, TBK^{+}20]. detectives [PCK^{+}13]. Developer [CWZF24]. Developing [HZKW21, KCTH19]. Development [FJZ^{+}22, MGP^{+}22, YK16]. Dialog [HBZ23, ZF17]. Dialogue [And12, MR19, RSvO^{+}22, SD23, ZT22]. Dialogue-based [SD23]. Diet [HMH^{+}19]. Differences [AYL^{+}22]. Different [CHST23]. Diffusion [AL20]. Digital [GRNW15, YDL^{+}17, YVA^{+}24]. Dimension [VPB16, WDN24]. Dimensional [GSC^{+}17, SDW^{+}18, EWS12]. Dimensionality [DM17]. Diminish [Hol18]. Direction [MKS^{+}17]. Directions [BPUK18, JJ17, NBHJ16]. Directive [SML^{+}23]. Disclosure [KK13, SW20]. Discourse [RSvO^{+}22]. Discover [JLH^{+}23]. Discovering [CMPS17, RD21, ZF17]. Discovery [KB16, SGR^{+}20]. Discussions [RK16]. Disease [MBD^{+}15]. Disorders [MDR^{+}22, WZB^{+}16]. Display [NWP^{+}18]. displayed [BSBC12]. displays [MGF13]. Distancing [DM17]. Distinguished [Afe18]. Distraction [DPR^{+}22]. Diverse [BHNP20, PCNB17]. Diversity [KB17, SFKL20, TB20]. Diversity-Promoting [TB20]. Do [GRC^{+}12, SF23, SW20]. Document [SMM20]. Does [MNL^{+}23]. Domain [DMGG13, ITB19, OHZ^{+}18, OLS21, SG12].

Domain-Independent [ITB19].

Dominance [NYYT16]. don’t [EBB^{+}22]. Drama [CSD14]. Drawing [DHS^{+}17]. Drift [MKS^{+}17]. Driven [DSV17, HBKM20, PBKS20, GCDN11, KPP14, SFTI^{+}13]. Driver [DPR^{+}22]. Drivers [HWM^{+}22, KK12]. Driving [HWM^{+}22, WZB^{+}16]. Drone [TRO^{+}19]. During [SLMK18, GSC^{+}17, YLVE12]. dyadic [HCDRB12]. Dynamic [OHZ^{+}18, TVCL16]. Dynamics [SKG^{+}18]. Dyslexic [WCS20].

E-Commerce [BHNP20]. E-learning [WCS20]. Eating [LPBG19]. EDAssistant [LZL^{+}23]. Editing [TVCL16]. Editorial [LC18, Zho22]. Educational [WBL^{+}13]. EEG [AHJR18]. Effect [BLIC21, CLS^{+}18, HHC20, KKL^{+}19, LDM23, ZLW17, ZMLY19]. Effective [RHS19, KR14]. Effector [CMGM18]. Effects [CRO^{+}16, CHST23, DM17, HBZ23, RFR^{+}24, SHY16, WY22]. Efficient [FLJA23, LBC^{+}21, NLN^{+}14, SWS21]. Elicit [EVB^{+}20]. Eliciting [HCDRB12]. Email [HHC20]. Embedded [TTL^{+}22]. Embeddings [AL22]. Embodied [OPM17, PAGM15, ZLW17]. Embodying [TDBL13]. Emerging [SCLH21]. emoji [SKS^{+}24]. Emojis [KCTH19]. Emotion [BEC^{+}20, MKW21, PSOC16, WES16]. Emotional [BSBC12, DLMC15, PW19, SP22, MGP^{+}22]. emotionally [DG12]. Empathy [BEC^{+}20, PLBW17]. Empirical [BPUP22, CGT12, KB17, KDA^{+}22]. Employing [CQW^{+}14]. Enable [SG12]. Enabled [BDO^{+}16, MDR^{+}22]. Enabling [FLJA23]. Encoding [VSR12]. Encounters [FLT20]. End [CMGM18, FJZ^{+}22, NSA^{+}22, KSW^{+}11]. End-Effector [CMGM18]. End-to-End [FJZ^{+}22]. end-user [KSW^{+}11]. End-users [NSA^{+}22]. Ended [YLC23]. Energy [SWS21]. Energy-Efficient [SWS21]. engagement [INN13]. Engaging [HCD^{+}20]. English [DMGG13]. English-Language [DMGG13]. Enhance [CMPS17]. Enhanced [WCS20, ZT22]. Enhancement [SSL^{+}19]. Enhancing [KCC^{+}19]. Ensemble [BHNP20]. Entertainment [SHA^{+}15b, SHA^{+}15a]. Enthusiasts [PDA^{+}21]. Entity [YVA^{+}24]. Environment [DFR15, LCKM23, SHY16, VLP^{+}15, WZB^{+}16, TDBL13]. Environments [AG13, MBD^{+}15, TSSL15, YLC23, CRP^{+}12]. Errors [KKL^{+}19]. Estimating [INN13, NWP^{+}18]. Estimation [APY^{+}20, CMTB15, NYYT16]. Ethics [Shn20, VPAJ^{+}23]. Evaluating [EBB^{+}22]. Evaluation [GCDN11, HC16, KCM22, MNL^{+}23, MKS^{+}17, MZR21, Riv14, ZCC^{+}18, YIS^{+}14]. Evaluations [CYW19]. Event [DPS^{+}19, KP18, MSD^{+}16, PCK^{+}18]. EventAction [DPS^{+}19]. Events [SB21]. Everyday [EVB^{+}20]. Examining [SF23]. Examples [SLVV23]. Exercise [MVH20]. ExG [SCLH21]. ExG-based [SCLH21]. Experience [JIKB22, SF23, KK12]. Experiences [COAR16, MDR^{+}22]. Experimental [BHNP20, MNL^{+}23]. Experiments [CSD14]. Expert [GGK^{+}21, HKR^{+}18, LDM23]. Expertise [MTML20, OHZ^{+}18, OLS21]. Explain [KLC^{+}21, ZFM^{+}24]. Explainability [SML^{+}23, SKS^{+}24]. Explainable [AF22, DNV^{+}23, DPS^{+}19, MZR21, PBF^{+}23, RNA^{+}23, TVA^{+}21, VPAJ^{+}23]. Explaining [DDGMP20, HBZ23]. Explanation [DMGG13, KLC^{+}21, LDM23, MNL^{+}23, SJS^{+}21]. Explanations [ADS^{+}20, CHST23, CYW19, KSP^{+}20, LDM23, LDLW23, MJJT24, PDA^{+}21, SML^{+}23, SV22, WY22]. Explicit [JWK14]. Exploitation [BHNP20]. Exploiting [RAMZ^{+}17]. Exploration [AUMT16, BHNP20, CCM14, GGC^{+}19, MSD^{+}16, PCK^{+}18, SGR^{+}20, ZGZ^{+}16]. Exploration/Exploitation [BHNP20]. Exploratory [AG13, DSV17, GZGW23, GSG^{+}24, KMO22, LZL^{+}23, DBTV20, SKKM15, TRO^{+}19]. Explorer [HSS^{+}21]. Exploring [FWY^{+}18, HSS^{+}21, HC16, Moh21, SKG^{+}18, TSSL15, TB20, VBCC20, YDL^{+}17, JJT24]. Expression [MKS^{+}17]. Expressions [BEC^{+}20, TVCL16]. Expressive [AF22, CMGM18, RPRK^{+}21]. Extended [CLJ21, SCLH21]. Extending [ZGZ^{+}16]. External [KW16]. Extraction [DAW18, LDS^{+}22, SKKM15]. Eye [CLRT20, MNL^{+}23, NBHJ16, SLH^{+}21, SCC14, XZY16, AC13, INN13]. Eye-tracking [CLRT20, MNL^{+}23]. Eyewear [MKS^{+}17].

Face [MBR16, MTML20]. Faceted [GZGW23]. Facial [BEC^{+}20, MKS^{+}17]. Facilitating [SGR^{+}20]. Factorisation [AF22]. Factorisation-based [AF22]. factorization [RAE^{+}14]. Fairness [NSA^{+}22]. False [VOA^{+}22]. FAtiMA [MGP^{+}22]. Faults [KDA^{+}22]. Feature [AF22]. Features [CMPS17, OHZ^{+}18, PSMO18]. Federated [MWP^{+}21, Yan21]. Feedback [FB18, JWK14, NLN^{+}14]. Few [SMF^{+}23, SLVV23]. Field [VPAJ^{+}23]. Finder [YNK^{+}22]. Finding [CCM14, KDA^{+}22, YNK^{+}22]. five [EWS12]. five-dimensional [EWS12]. Floor [VLP^{+}15]. Fluid [ZBGD14]. Focus [CDK^{+}18]. Follow [ITB19]. Follow-Up [ITB19]. Footprint [LBC^{+}21]. Footprinting [YVA^{+}24]. Foraging [PDA^{+}21]. ForSense [RRS^{+}22]. Forums [FWY^{+}18]. Fostering [MJJT24]. FourEyes [SFKL20]. Framework [BHNP20, BEH22, CSO^{+}22, DPA20, GKP15, HRAR18, JWK14, LL23, MZR21, PKCS12, PSMO18, SKG^{+}18, WRH18]. free [KPMHB22, YNK^{+}22]. Freezing [MBD^{+}15]. Full [NKP^{+}19]. Full-Body [NKP^{+}19]. functionality [KPP14]. Fusing [BFOK12]. Future [BPUK18].

Gait [MBD^{+}15]. Galvanic [NCWC17]. Game [CLP^{+}14, KCM22, LCKM23, PDA^{+}21, TSSL15, YLC23]. Game-Based [TSSL15]. Gameplay [LCKM23]. Games [SHKM19]. Gamification [SJS^{+}21]. Gap [Shn20]. gastronomy [CAB^{+}13]. Gaze [BLIC21, INN13, IOKY16a, JFNY13, NYYT16, NBHJ16, PAGM15, RAMZ^{+}17, SLH^{+}21, SCC14, WZB^{+}16, YYI^{+}16, AC13]. Gaze-Contingent [WZB^{+}16]. Gaze-Exploiting [RAMZ^{+}17]. generAItor [SKS^{+}24]. General [VLP^{+}15]. General-Purpose [VLP^{+}15]. Generalisable [SD23]. Generate [KLC^{+}21]. generated [DSEV19]. Generating [KSP^{+}20, NYYT16, SV22, TTL^{+}22]. Generation [AMN^{+}13, CMGM18, CWZF24, DMGG13, FJZ^{+}22, SMF^{+}23, SKS^{+}24, YN13]. Generative [KLC^{+}21]. Genome [VSR12]. Geometrically [RDP^{+}24]. gestural [ZBGD14]. Gesture [CMTB15, KKL^{+}19, KCTH19, PSOC16, TSSL15, SDD12, ZBGD14]. Gestures [AUMT16, KCTH19, MBR16, SLVV23, YYI^{+}16]. Givers [TOF16]. GO [YNK^{+}22]. GO-Finder [YNK^{+}22]. GRAFS [GZGW23]. Graphical [GZGW23, KTD17, TJLO20, VBCC20]. Ground [DAW18]. Group [AYL^{+}22, CRO^{+}16, EDR24]. Groups [CSD14, FWY^{+}18, KR14]. GSR [AHJR18]. Guest [LC18]. GUI [BLO22]. Guidance [BLIC21]. Guidelines [Shn20].

Habitual [RD21]. Hand [KCTH19, MBR16, OLS21, SLVV23, YNK^{+}22, SDD12].

Hand-held [YNK^{+}22]. Hand-Over-Face [MBR16]. Handwriting [OHZ^{+}18]. HCI [DDGMP20]. Health [BPUP22, EBB^{+}22, KMO22, MVH20, Moh21]. Healthcare [KLC^{+}22, HBP^{+}12]. held [YNK^{+}22]. Help [MNL^{+}23]. Helpers [COAR16]. Heterogeneous [AL23]. Hierarchical [CKKD14]. High [GSC^{+}17, MSD^{+}16, SDW^{+}18]. High-Dimensional [GSC^{+}17, SDW^{+}18]. High-Volume [MSD^{+}16]. Highlights [BBQ20, BCKC20, CDF19, HKOT22, JR12]. HILC [ITB19]. Historical [HMO22, MTML20]. History [HK16]. Holistic [MLD^{+}21]. Home [DNV^{+}23, DFR15]. Homeless [JIKB22]. Household [HBKM20]. Human [AHJR18, And12, BPUK18, BMD^{+}15, BHS^{+}21, CSO^{+}22, CdGF^{+}13, DM17, FG18, Gil15, HMH^{+}19, JIKB22, JJT24, JCL^{+}19, KFWL14, KP18, KGCC18, KCM22, LC18, MR19, MTML20, NSA^{+}22, PBKS20, PBF^{+}23, RFR^{+}24, RAMZ^{+}17, RK16, Shn20, TSN^{+}16, TOF16, VLP^{+}15, WRH18, WHLZ18, XZY16, Zho21, GR13, HCDRB12, ONTHS11, SDD12, ZFM^{+}24]. Human-Agent [TSN^{+}16, JIKB22]. Human-AI [CSO^{+}22]. Human-Centered [FG18, Zho21, BHS^{+}21, PBF^{+}23, Shn20]. Human-Computer [And12, PBKS20, SDD12]. Human-Environment [VLP^{+}15]. Human-in-the-Loop [KP18, WHLZ18, HMH^{+}19, KCM22, NSA^{+}22]. Human-Machine [JCL^{+}19, KGCC18]. Human-Robot [BPUK18, WRH18, XZY16, HCDRB12, ONTHS11]. Humanized [TFT21]. Humanoid [CNI15]. Humans [RSPT^{+}23, WWB^{+}23]. Hybrid [CSO^{+}22, KSP^{+}20]. Hypothesis [MSD^{+}16].

Icon [FJZ^{+}22, SFTI^{+}13, FJZ^{+}22]. icon-driven [SFTI^{+}13]. idea [YN13]. Identification [LPBG19, SVK16].

Identifying [MTML20]. I’ll [YYI^{+}16]. Illmo [Mar14]. Illocutionary [SV22]. Image [AGH^{+}23, TBK^{+}20]. Images [DDGMP20]. Immersive [RSPT^{+}23]. Impact [WSL^{+}18, YLC23]. Impactful [HHC20]. Impairments [FLJA23, AFFB14, CSM^{+}12]. Implicit [JWK14]. Importance [NRB^{+}22]. Impressions [NRB^{+}22]. Improve [GGK^{+}21, SFKL20]. Improving [KKS22]. In-the-Wild [MDR^{+}22]. Inclusive [GRNW15]. Incremental [DFR15]. Independent [ITB19]. indicators [CRC^{+}12]. Individualising [TJLO20]. Individually [BFOK12]. Individuals [WZB^{+}16, CSM^{+}12]. Inducing [CSD14]. Inference [SCKOGP15]. Inferring [KW16, RAMZ^{+}17, SCC14, WDN24]. Influence [KGCC18, YC23]. Influencing [BFOK12]. Information [BPUP22, CMPS17, CLJ21, GGC^{+}19, KK13, KLR^{+}18, SF23, SLH^{+}21, SW20]. Initial [VOA^{+}22]. initiated [ZCC21]. Input [KKL^{+}19, vdGPOvB^{+}13]. input-output [vdGPOvB^{+}13]. Inputs [EBB^{+}22, SCLH21]. Inspection [AYL^{+}22]. Installation [GRNW15]. Instruction [Gil15]. Instructions [BLV14]. Instrumented [CSD14]. Insulin [SF23]. Integrated [AFFB14, DSEV19]. Integration [RRS^{+}22]. Integrity [MJJT24]. Integrity-based [MJJT24]. Intelligence [GGK^{+}21, Zho21, PCK^{+}13]. Intelligent [BBQ20, BCKC20, BEH22, CGHS17, CDF19, EVB^{+}20, EGQ16, GRC^{+}12, JR11, KW16, KGCC18, NBHJ16, PBKS20, PASR^{+}16, YLC23, ZCC21, CAB^{+}13, DG12, JR12, SFTI^{+}13]. Intent [KLR^{+}18, WDN24]. Interacting [CAB^{+}13, JJ17, RSvO^{+}22]. Interaction [BPUK18, BBT^{+}15, CLRT20, CNI15, EGQ16, FLJA23, FB18, KFWL14, KGCC18, KB16, LL23, NYYT16, PBKS20, SB21, SDW^{+}18, TSN^{+}16, TOF16, VLP^{+}15, VSR12, XZY16, YC23, ZLW17, ZCC21, CRP^{+}12, CRC^{+}12, ONTHS11, SLM^{+}13, SDD12, ZBGD14, vdGPOvB^{+}13]. Interactions [BMD^{+}15, DGG16, DM17, JCL^{+}19, KCC^{+}19, Moh21, VOA^{+}22, WDN24, ZGZ^{+}16, HCDRB12, JFNY13]. Interactive [ABJ15, BSKB21, CMTB15, CGHS17, CCM14, CFD^{+}14, CKKD14, DSEV19, DK18, GRNW15, HMO22, HC16, JR11, KTD17, KGCC18, LH12, LC18, LDS^{+}22, Mar14, MKF^{+}15, NBHJ16, NSA^{+}22, NLN^{+}14, NRB^{+}22, NSMM18, PCNB17, PCK^{+}18, RFR^{+}24, SVK16, SJS^{+}21, SMM20, SKKM15, SGR^{+}20, SG12, TVA^{+}21, ZCC21, GCDN11, HJ11, JR12, KPP14, YIS^{+}14]. Interactivity [HBZ23]. Interest [LBC^{+}21]. Interface [AL22, AL23, BBQ20, BCKC20, CDF19, DMGG13, DK18, GSG^{+}24, HBZ23, SWS21, TVCL16, SFTI^{+}13]. Interfaces [BSKC21, FLJA23, SV22, TB20, YC23, ZTD^{+}24, ZBGD14]. Interference [NCWC17]. Intergroup [Hol18]. Internal [RHS19]. Internet [GR13, PCK^{+}13, YN13]. Internet-scale [GR13, PCK^{+}13, YN13]. Interpersonal [CRO^{+}16, JCL^{+}19]. Interpretations [PSMO18]. Interpreting [AL20, BLV14, RDP^{+}24]. Intervention [RHS19, WZB^{+}16]. Interventions [YLC23]. Introduction [AC13, BCKC20, BBT^{+}15, CRP^{+}12, CGHS17, CFD^{+}14, FG18, GR13, HKOT22, JR11, JR12, LH12, NBHJ16, SHA^{+}15b, SLM^{+}13, TOF16, Zho20, Zho21, Zho22]. Investigating [CGT12]. Investigation [HMO22]. Investigations [HZKW21]. Involvement [BEH22]. Involving [NSA^{+}22]. IoT [HBKM20]. Isolated [SBN^{+}18]. Issue [BBQ20, BCKC20, BBT^{+}15, CGHS17, CDF19, CFD^{+}14, FG18, HKOT22, KGCC18, KLC^{+}22, LH12, LC18, NBHJ16, PBKS20, SHA^{+}15b, TOF16, TVA^{+}21, CRP^{+}12, JR12, SLM^{+}13]. Issues [TFT21]. item [RAE^{+}14]. Items [SHK19]. IUI [BBQ20, BCKC20, CDF19, HKOT22].

John [Jam13]. Just [HKR^{+}18].

kernel [HJ11]. Key [SP22]. Keyterm [NSMM18]. Keyterm-Based [NSMM18]. Know [OLS21, ZFM^{+}24, EBB^{+}22]. Knowledge [BEH22, GRC^{+}12, KH12, VSR12].

Lab [MBD^{+}15]. Labeling [BHS^{+}21, RFR^{+}24, SJS^{+}21]. Language [BLV14, DMGG13, FBC14, GSG^{+}24, PLBG19, SKS^{+}24, BSBC12, PCK^{+}13]. Large [BSKB21]. Latent [AF22]. laughter [SGS^{+}12]. Layouts [TJLO20]. Leading [DGG16]. Learn [KLC^{+}21]. Learned [JCL^{+}19, NRB^{+}22]. Learner [BDO^{+}16]. Learning [AG13, ADS^{+}20, AL22, AL23, BHS^{+}21, BLO22, CDK^{+}18, CFD^{+}14, CKKD14, DFR15, DG12, DK18, FG18, FB18, HMH^{+}19, HTOB^{+}20, JJL16, KFWL14, KTD17, KLC^{+}21, KCC^{+}19, KPSK17, MWP^{+}21, MHC^{+}18, NLN^{+}14, PYDM16, RHS19, RNA^{+}23, SHK19, SML^{+}23, SD23, SGR^{+}20, WCS20, Yan21, ZTD^{+}24, ZLG22, ZTSH23, HJ11]. Lessons [NRB^{+}22]. Level [KKL^{+}19, SDW^{+}18]. Levels [CHST23, LCKM23]. Leveraging [SFKL20]. like [HCDRB12]. LIMEADE [LDLW23]. linear [BFY^{+}21]. lingual [SFTI^{+}13]. Linguistic [SJS^{+}21]. Listening [YT15]. Lists [GRC^{+}12]. Literacy [BLIC21]. LiveAction [AMN^{+}13]. Living [MDR^{+}22, SD23]. Load [NCWC17, CRC^{+}12]. Local [MNL^{+}23]. localization [AFFB14]. Localized [FBC14, PSMO18]. Logic [CHST23]. Logic-Style [CHST23]. Long [ZT22]. Long-term [ZT22]. Looks [ZFM^{+}24]. Loop [KP18, WHLZ18, HMH^{+}19, KCM22, NSA^{+}22, SKS^{+}24]. Lost [YNK^{+}22, ZF17].

Machine [CDK^{+}18, CFD^{+}14, DK18, FG18, FB18, JJL16, JCL^{+}19, KFWL14, KTD17, KLC^{+}21, KGCC18, MHC^{+}18, RFR^{+}24, RRS^{+}22, SML^{+}23, SGR^{+}20, ZCC21, ZTD^{+}24, ZTSH23]. Machine-initiated [ZCC21]. Machine-Learning [KTD17]. Machines [AHJR18]. Magazine [BLIC21]. Magazine-Style [BLIC21]. Major [WWB^{+}23]. Making [CdGF^{+}13, DM21, GGK^{+}21, KTD17, KK13, SLMK18, TVA^{+}21, WY22, HSS^{+}21]. Man [RFR^{+}24]. Management [PAGM15]. Manipulating [LZW^{+}24]. Manipulation [KPMHB22]. Mapping [FB18, KCTH19]. Maritime [Riv14]. Markov [DMGG13]. Math [TVCL16]. Mathematical [TVCL16]. Matrix [AF22, CLJ21]. Matter [HKR^{+}18]. Me [XZY16, CWZF24]. mean [EBB^{+}22]. Meaningful [LDM23]. Means [SFKL20]. Measured [DM21]. Measures [BHS^{+}21, MHC^{+}18]. Mechanism [LL23]. Media [PKCS12]. Mediated [KCTH19]. Medical [DAW18, LDM23]. Meet [WWB^{+}23]. Meetings [IOKY16a, IOKY16b]. Memoriam [Jam13]. Mental [ADS^{+}20, EBB^{+}22, KMO22]. Menus [VBCC20]. Mere [ADS^{+}20]. Messages [KDMA12, KPMHB22]. Metaphors [ABJ15]. Method [EBB^{+}22, EVB^{+}20]. Methodology [GCDN11]. Methods [CLJ21, GSC^{+}17, MKF^{+}15]. Metrics [BSKC21, HKR^{+}18]. Metrics-based [BSKC21]. MI3 [ZCC21]. Micro [LL23]. Micro-moment [LL23]. Mid [AUMT16, SLVV23]. Mid-Air [AUMT16, SLVV23]. Mindsets [BEH22]. Minimal [KB16]. Mining [YSD15]. Miscommunication [MR19]. Mitigating [RD21]. Mitigation [DJH^{+}12, PKCS12, RDP^{+}24]. Mobile [AL22, BSKC21, CSD14, HMH^{+}19, PW19, MGF13, TDBL13]. Mobility [PSMO18]. MobInsight [PSMO18]. modal [AL22]. Modalities [CFD^{+}14]. Model [AHJR18, AMN^{+}13, BEC^{+}20, CQW^{+}14, HBZ23, HWM^{+}22, LZW^{+}24, MNL^{+}23, Moh21, Riv14, RSPT^{+}23, SVK16, SKS^{+}24, SP22, ZTD^{+}24, ZT22, KPP14]. Model-agnostic [MNL^{+}23]. model-driven [KPP14]. Modeling [GGK^{+}21, HKR^{+}18, HC16, JWK14, LPBG19, OBG^{+}19, PBKS20, WRH18, WCS20, YVA^{+}24]. Modelling [AF22]. Models [ADS^{+}20, FBC14, GGK^{+}21, RNA^{+}23, TJLO20, YPL^{+}16, YK16]. modern [ZBGD14]. moment [LL23]. Monitoring [HMH^{+}19, HRAR18, YVA^{+}24]. Mood [SCKOGP15]. Mortals [ADS^{+}20]. Motion [FB18, WHLZ18]. Motion-Sound [FB18]. Motivation [RHS19, WCS20]. Motor [MHC^{+}18]. Movement [ABJ15, NKP^{+}19]. Movements [CMGM18, OLS21, WES16]. Movie [KMAM20]. MovieLens [HK16]. Multi [AL22, BFY^{+}21, KCM22, MDC^{+}24, WHLZ18]. Multi-agent [KCM22]. Multi-modal [AL22]. Multi-Objective [MDC^{+}24, BFY^{+}21]. Multi-Robot [WHLZ18]. Multiclass [NLN^{+}14]. Multidimensional [CCM14]. Multidisciplinary [MZR21]. Multifaceted [WBL^{+}13]. Multifocus [PCK^{+}18]. Multilingual [ZT22]. Multimedia [PSC^{+}16]. Multimodal [AL23, CRC^{+}12, COAR16, DPR^{+}22, ONTHS11, PSC^{+}16, PW19, SCKOGP15, XZY16, CSM^{+}12]. Multiparty [IOKY16a, IOKY16b, KFWL14, NYYT16, SGS^{+}12]. Multiple [CFD^{+}14, KH12, MHC^{+}18, RSvO^{+}22, VPB16, HJ11]. Multiplex [YYI^{+}16]. multistroke [HP11]. multitask [EWS12]. Museum [CSD14, DK15]. Music [DLMC15, YT15, YLVE12].

naive [KSW^{+}11]. Narrative [BLIC21]. Narratives [SB21]. Natural [BLV14, GSG^{+}24, CRP^{+}12, EWS12, SGS^{+}12, SDD12]. Naturalistic [MBR16]. Nature [WBL^{+}13]. Navigation [BFY^{+}21, ZLW17, AFFB14]. Neighborhood [PSMO18]. Network [AL22]. Network-based [AL22]. Networks [AL23, LWFM23, CAB^{+}13]. Neural [LWFM23, ZT22]. Neuron [LWFM23]. News [HMO22]. Next [IOKY16a, IOKY16b, YYI^{+}16]. Non [BFY^{+}21, EGQ16, LDM23]. Non-Expert [LDM23]. Non-linear [BFY^{+}21]. Non-Visual [EGQ16]. Nonstrict [CKKD14]. Normal [Riv14, VBCC20]. Norms [SWS21]. Notebooks [LZL^{+}23]. Noteworthy [SB21]. Novel [VSR12]. Novelty [KB17]. NUI [SCLH21].

object [HJ11]. Objective

[MDC^{+}24, BFY^{+}21]. Objectives [KB17]. Objects [YNK^{+}22, MGF13, SLM^{+}13, vdGPOvB^{+}13]. Observation [SDW^{+}18]. Observation-Level [SDW^{+}18]. Occurrences [JLH^{+}23]. Office [KKS22]. offline [SGS^{+}12]. Older [BPUP22, RSvO^{+}22, SBN^{+}18]. Ongoing [ZF17]. Online [CWZF24, FWY^{+}18, HC16, PSC^{+}16, RRS^{+}22, SB21, AFFB14, KR14, SGS^{+}12]. Open [SG12, YLC23]. Open-Domain [SG12]. Open-Ended [YLC23]. Operations [MKF^{+}15]. Optimization [BFY^{+}21, MKF^{+}15, MDC^{+}24, MKW21, TTL^{+}22, ZTD^{+}24]. Order [YYI^{+}16]. oriented [HMO22, KSW^{+}11]. Out-of-the-Lab [MBD^{+}15]. Outcomes [BEH22, RHS19]. output [vdGPOvB^{+}13]. Overview [JJ17, Yan21]. Oz [BL15].

Paintings [SYS^{+}15]. Pairs [DK15]. Paper [ZB23]. Parametric [CQW^{+}14, SDW^{+}18]. Park [SKG^{+}18]. Parkinson [MBD^{+}15]. Participation [NYYT16]. partitioning [GGC^{+}19]. Path [HSS^{+}21]. Paths [HSS^{+}21]. Pattern [CCM14, SKKM15]. Patterns [BLIC21, HSS^{+}21]. PbD [ITB19]. PEACE [SP22]. Pedagogical [YLC23]. Pen [TVCL16]. Pen-Based [TVCL16].

People [FLJA23, HBP^{+}12, MDR^{+}22, AFFB14, CAB^{+}13, KK12]. Perception [OPM17]. Perceptions [Hol18]. Perceptual [CMGM18]. Performance [SCC14, YLC23, YDL^{+}17]. Performing [YDL^{+}17]. Person [DPA20]. Person-Context [DPA20]. Personal [CGHS17, HTOB^{+}20]. Personality [And12, DPA20, KMAM20, PBKS20, SHKM19, TBK^{+}20, YC23, ZMLY19]. Personalization [BFOK12]. Personalized [CLJ21, HMH^{+}19, KSP^{+}20, LL23, MVT16, KK12, RAE^{+}14]. Personhood [Hol18]. Perspective [CGT12]. Persuasion [And12, BFOK12, CGT12]. Persuasive [KDMA12]. Persuasiveness [PSC^{+}16]. Phenomena [SJS^{+}21]. Phenomenon [WRH18]. Philosophy [SV22]. phones [AFFB14]. Photo [MKW21, MTML20]. Phrase [PCK^{+}13]. Physical [NKP^{+}19, TDBL13]. Physiological [TBK^{+}20, YLVE12]. picoTrans [SFTI^{+}13]. Pittsburgh [HCD^{+}20]. places [TDBL13]. Plan [AG13]. Planning [KH12, WHLZ18]. Players [PDA^{+}21]. Policy [AL20]. PolicyFlow [AL20]. Portraits [MTML20]. Position [APY^{+}20]. Positional [AL23, MKS^{+}17]. Positives [VOA^{+}22]. Possible [JJT24]. Potential [CGT12]. Powered [CWZF24]. Practice [MHC^{+}18, Shn20]. Practitioner [Zho22]. Predict [IOKY16b, OHZ^{+}18]. Predicting [EDR24, KMAM20, SLMK18, SLH^{+}21, WBL^{+}13]. Prediction [CLRT20, DLMC15, FBC14, IOKY16a, PSC^{+}16, RK16]. Predictions [ZTD^{+}24]. Predictive [TJLO20]. Preference [DM17, FBS13, KMAM20]. Preferences [HTOB^{+}20, JWK14]. Presence [CRO^{+}16]. Presentation [HHC20]. Presentations [ML19]. preserving [Yan21]. Prevention [DJH^{+}12]. Primary [KLR^{+}18]. PRIME [CLJ21]. primitives [HP11]. Principles [WY22, KK12]. Privacy [HBKM20, KK13, Yan21]. Privacy-preserving [Yan21]. Proactive [KLR^{+}18]. Probabilistic [RNA^{+}23]. Problem [FLT20, GR13]. Problems [EVB^{+}20]. Process [YK16]. Processes [DMGG13]. Processing [ML19, SF23]. Professional [SYS^{+}15]. Profiles [EDR24]. Profiling [SHKM19]. Progressive [SW20]. Projected [HSS^{+}21, MGF13]. Projection [HSS^{+}21]. Projective [RSPT^{+}23]. Promoting [MVH20, SWS21, TB20]. Proof [RSPT^{+}23]. Proof-of-concept [RSPT^{+}23]. propagation [RAE^{+}14]. Properties [SCC14]. Property [BHS^{+}21]. Provenance [CQW^{+}14]. Providing [JIKB22, RK16]. Proxemics [CRO^{+}16]. Proxy [FLJA23]. Psychological [RSPT^{+}23]. Public [ML19, NWP^{+}18]. Purpose [VLP^{+}15].

Qualitative [CDK^{+}18]. Qualities

[ABJ15, SP22]. Quality

[CGT12, DSEV19, GGC^{+}19, HCD^{+}20, NKP^{+}19, WBL^{+}13, ZTD^{+}24]. Quality-assuring [ZTD^{+}24]. Quantify [PW19]. Quantifying [DHS^{+}17]. Quantitative [YT15]. Quartet [DGG16]. Queries [SMF^{+}23]. Query [FBC14, HMO22]. Query-oriented [HMO22]. Question [GSG^{+}24, SV22]. QuestionComb [SJS^{+}21]. Questions [BEH22, ITB19].

Radar [APY^{+}20, SLVV23]. RadarSense [SLVV23]. Rank [KLC^{+}21, MLD^{+}21]. Rating [FBS13, KMAM20]. Raw [Gil15]. Read [EGQ16]. Real [BSKB21, CLRT20, PDA^{+}21, PASR^{+}16, RFR^{+}24, WES16]. Real-Time [PASR^{+}16, WES16, BSKB21, CLRT20, PDA^{+}21, RFR^{+}24]. Realities [SCLH21]. Reality [WZB^{+}16]. realization [KPP14]. Reasoning [DJH^{+}12, KH12, SG12]. Recognition [AG13, APY^{+}20, BL15, BBT^{+}15, CMTB15, CLP^{+}14, DPR^{+}22, DNV^{+}23, DPA20, KCTH19, MKS^{+}17, MTML20, PLBG19, PSOC16, PASR^{+}16, RNA^{+}23, SVK16, SLVV23, TSSL15, WES16, YSD15, YK16, SDD12, YLVE12, ZBGD14]. Recognizer [BL15]. Recognizing [CNI15, HP11]. Recommendation [BHNP20, CYW19, DLMC15, DPS^{+}19, HTOB^{+}20, JJL16, LZL^{+}23, MLD^{+}21, PYDM16, VPB16, WWB^{+}23, YC23, RAE^{+}14]. Recommendations [DDGMP20, HBZ23, PCNB17, TB20]. Recommender [AF22, CdGF^{+}13, CLJ21, CGT12, JWK14, KB17, KK13, KSP^{+}20, KB16, LL23, SHK19, SWS21, TFT21]. Recommenders [JJ17]. Recommenders-Overview [JJ17]. Recommending [MVT16]. Reconstruction [ZCC21]. Record [FLJA23]. Recovery [MR19]. Reduce [KDMA12]. Reduction [WDN24]. Refinement [ZCC^{+}18]. Reflections [MHC^{+}18]. Registration [YNK^{+}22]. Registration-free [YNK^{+}22]. Regression [CCM14]. Regular [MVH20]. Reinforcement [ADS^{+}20, CKKD14, ZLG22]. Relation [DAW18]. Relational [JIKB22]. relations [vdGPOvB^{+}13]. Relationships [JIKB22]. Relevance [VPB16]. Relevant [KPMHB22]. Reliability [CLS^{+}18]. Reliable [Shn20]. Repair [BPUK18]. Repeatability [MKS^{+}17]. Repeated [KKL^{+}19]. report [MDR^{+}22]. Reported [EVB^{+}20]. Representation [PSOC16, RAMZ^{+}17]. Representations [MNL^{+}23, RDP^{+}24]. Requesting [SHY16]. Research [BPUK18, CSO^{+}22, JJ17, MKF^{+}15, RRS^{+}22, RSPT^{+}23, TFT21, VPAJ^{+}23, VPB16, WSL^{+}18]. resource [PCK^{+}13]. Resources [WBL^{+}13]. Respiration [IOKY16b]. Response [GKP15, NCWC17, YDL^{+}17]. Responses [EBB^{+}22, PW19, TBK^{+}20, VOA^{+}22]. Responsibility [DM21]. Responsible [Yan21]. Responsive [BEC^{+}20]. Results [ZCC^{+}18]. Retrieval [KLR^{+}18, HJ11]. Reuse [LDS^{+}22]. Reveal [OLS21]. Review [DKI^{+}21, DK18, MKF^{+}15]. Reviewers [Afe18]. RF [APY^{+}20]. RF-Radar [APY^{+}20]. Rich [AL22]. Riedl [Jam13]. Risk [WRH18]. Roadmap [DBTV20]. Robot [BPUK18, CNI15, GKP15, KFWL14, MR19, NYYT16, RHS19, RPRK^{+}21, WRH18, WHLZ18, XZY16, YYI^{+}16, HCDRB12, ONTHS11]. Robotic [CLP^{+}14, Hol18]. Robots [CFD^{+}14, CKKD14, PLBW17, RAMZ^{+}17]. Role [Moh21, VPAJ^{+}23, XZY16]. Roles [NYYT16]. Routines [DFR15].

Safe [Shn20]. Say [SG12]. Scalable

[PCNB17]. scale [GR13, PCK^{+}13, YN13]. Scattered [ZF17]. Scene [LDS^{+}22]. Science [CDK^{+}18, HCD^{+}20]. Sclerosis [MHC^{+}18]. Screen [KPMHB22, KKS22]. Screen-free [KPMHB22]. Screens [VBCC20]. scrutinize [KK12]. Search [CMPS17, DSV17, GZGW23, KLR^{+}18, LZL^{+}23, DBTV20, SLH^{+}21, TJLO20, ZCC^{+}18]. section [AC13, GR13]. See [XZY16]. Seeking [BPUP22]. Selection [GSC^{+}17]. Self [KKS22, MDR^{+}22]. Self-adjusting [KKS22]. Self-report [MDR^{+}22]. Semantic [PSMO18, RAMZ^{+}17, SSL^{+}19, SGR^{+}20, YSD15]. Semantically [AL22]. Semantics [AL23]. Semi [SKKM15]. Semi-Structured [SKKM15]. Semisupervised [PYDM16]. Sense [DJH^{+}12, KH12, LH12]. Sensed [MHC^{+}18]. Sensemaking [RRS^{+}22]. Sensing [AHJR18, APY^{+}20, SLVV23, TBK^{+}20, VLP^{+}15, EWS12]. Sensitive [CLP^{+}14, SMF^{+}23]. Sensor [KTD17]. sensors [HBP^{+}12]. Sentiment [CYW19]. Sentiment-based [CYW19]. Sequence [MSD^{+}16]. Sequences [DPS^{+}19, MKW21, PCK^{+}18]. Sequential [BLIC21]. Serendipity [KB17]. Series [KTD17, LPBG19]. Service [OBG^{+}19, YYI^{+}16]. services [KPP14, TDBL13]. Sessions [FLT20]. Sets [SHK19]. Shape [CWZF24]. shapes [LBC^{+}21]. Shared [LCKM23]. Shifting [CDK^{+}18]. Should [JJT24, ZFM^{+}24]. Shoutcasters [PDA^{+}21]. sight [CSM^{+}12]. Sign [PLBG19]. Signal [ML19, OHZ^{+}18]. Signals [DPR^{+}22]. Signatures [HHC20]. SignCom [GCDN11]. signers [GCDN11]. Simulating [RSPT^{+}23]. Simulation [ZTD^{+}24]. Simulation-based [ZTD^{+}24]. Situ [LZL^{+}23]. Situated [MR19]. Situation [KPMHB22]. Situation-Relevant [KPMHB22]. Sketch [BL15, LDS^{+}22]. sketched [HP11]. Sketching [HKR^{+}18]. SketchMaker [LDS^{+}22]. Skills [TSN^{+}16]. Skin [NCWC17]. Sleep [HRAR18]. Sleuth [MTML20]. Small [CSD14, CNI15, VBCC20]. Smart [DNV^{+}23, KPMHB22, MKS^{+}17, AFFB14, MGF13, SLM^{+}13, vdGPOvB^{+}13, DFR15]. Smarter [ML19]. Smartphone [RD21, YT15]. Smartphones [ZGZ^{+}16]. SmartShots [TTL^{+}22]. Smell [HCD^{+}20]. Smiling [OPM17]. Snacking [KDMA12]. Social [BMD^{+}15, CDK^{+}18, CNI15, DK15, JIKB22, KFWL14, ML19, PSC^{+}16, RPRK^{+}21, SCKOGP15, DBTV20, SWS21, SP22, TSN^{+}16, TB20, CAB^{+}13, KR14, RAE^{+}14]. Socially [BFY^{+}21]. Socio [MGP^{+}22]. Socio-emotional [MGP^{+}22]. Software [CWZF24]. solving [GR13]. Sound [FB18, KP18]. Space [VBCC20]. Spatial [LZW^{+}24, ZLW17]. Spatialization [SSL^{+}19]. Speak [IOKY16b]. Speaker [IOKY16a]. Special [BBQ20, BCKC20, BBT^{+}15, CDF19, CFD^{+}14, FG18, HKOT22, KGCC18, KLC^{+}22, LC18, NBHJ16, PBKS20, SHA^{+}15b, TOF16, TVA^{+}21, Zho20, Zho21, Zho22, AC13, CRP^{+}12, GR13, JR12, SLM^{+}13, CGHS17, LH12].

Spectrum [WZB^{+}16]. Speech

[MDR^{+}22, EWS12]. Speech-enabled [MDR^{+}22]. Speed [DM17]. Spotting [SGS^{+}12]. Stability [YPL^{+}16]. Stance [KPSK17]. Standardized [EBB^{+}22]. State [TFT21]. State-of-the-art [TFT21]. States [DLMC15, YVA^{+}24]. Static [LZW^{+}24]. Statistical [SYS^{+}15, YPL^{+}16]. Statistics [Mar14]. Stimuli [TBK^{+}20]. Stimulus [GKP15]. Stimulus-Response [GKP15]. Storytelling [SG12]. Strangers [GRNW15]. Strategy [PDA^{+}21]. Streams [RFR^{+}24]. String [DGG16]. Structural [BSKC21]. Structure [ZT22]. Structured [JLH^{+}23, SKKM15]. Student [WCS20, YLC23]. Study [BPUP22, BHNP20, CGT12, KDMA12, KDA^{+}22, KLC^{+}21, KMO22, MDR^{+}22, VOA^{+}22, WWB^{+}23, YT15]. Studying [MVH20]. Style [BLIC21, CHST23]. stylistic [YIS^{+}14]. Subjective [DM21, SF23]. Subspaces [JLH^{+}23]. Success [SLH^{+}21]. Support [BEH22, CDK^{+}18, GZGW23, JIKB22, KMO22, LBC^{+}21, PBF^{+}23, RRS^{+}22, SBN^{+}18, VSR12, KR14]. Supported [VOA^{+}22]. Supporting [DSV17, JJL16, KPMHB22, LZL^{+}23]. Surfaces [APY^{+}20]. Survey [AGH^{+}23, KB17, MZR21, PLBW17, SCLH21]. Symbolic [KCTH19, WHLZ18, YK16]. Synthesizing [LCKM23]. System [And12, BL15, CLJ21, EGQ16, FWY^{+}18, GZGW23, ITB19, JJL16, KP18, KCTH19, KCM22, LDM23, YNK^{+}22, YYI^{+}16, YDL^{+}17, GCDN11, YN13]. Systematic [MSD^{+}16, YK16]. Systems [AF22, BEH22, CMTB15, CdGF^{+}13, CGT12, CFD^{+}14, CKKD14, DNV^{+}23, GGK^{+}21, JWK14, KB17, KDMA12, KK13, KGCC18, KSP^{+}20, KB16, LL23, MZR21, NBHJ16, SF23, SHK19, Shn20, SLH^{+}21, TFT21, WCS20, HJ11, JR12, KK12, CGHS17, JR11, LH12].

Tag [VSR12, RAE^{+}14]. tagging [RAE^{+}14]. Tailored [KDMA12]. Taking [LDLW23, JFNY13]. talk [DG12, GSG^{+}24]. talking [DG12]. Talks [VPB16]. Taming [GGK^{+}21]. tangible [vdGPOvB^{+}13]. Task [AMN^{+}13, KLR^{+}18, KPP14, PASR^{+}16, SLH^{+}21, SCC14]. Tasks [CLRT20, GRC^{+}12]. Taxonomy [BHS^{+}21, MKF^{+}15]. Teaching [TSN^{+}16]. Teaming [CSO^{+}22]. Teammates [JJT24]. Teams [ZFM^{+}24]. Technique [CCM14, YSD15]. Techniques [PLBG19, RDP^{+}24, TRO^{+}19, YIS^{+}14]. Technologies [SBN^{+}18, HBP^{+}12]. Technology [MNL^{+}23]. Teleoperation [TRO^{+}19]. tensor [RAE^{+}14]. term [ZT22]. Testing [MSD^{+}16]. Text [HMO22, KDMA12, SKS^{+}24, KSW^{+}11]. Textflow [KPMHB22]. Textual [SG12]. Their [KW16]. Theme [SKG^{+}18]. Theoretical [DM21]. Theory [GGC^{+}19, JCL^{+}19, SYS^{+}15]. There [YYI^{+}16]. things [CAB^{+}13]. Think [FLT20]. Think-aloud [FLT20]. Thread [ZF17]. Thresholding [KKL^{+}19]. Thumbnail [ACMH23]. TiiS [KGCC18, Zho20, Zho22, ZB23]. Time [DM17, GGC^{+}19, KTD17, LPBG19, PASR^{+}16, WES16, BSKB21, CLRT20, PDA^{+}21, RFR^{+}24]. Time-correlation-partitioning [GGC^{+}19]. Time-Series [KTD17]. To-Do [GRC^{+}12]. ToM [RSPT^{+}23]. too [CWZF24]. Tool [BSKC21, FJZ^{+}22, MGP^{+}22, SFKL20, ZTSH23]. Toolbox [Zho22]. Toolkit [DSEV19, MGP^{+}22]. Tools [CWZF24]. Top [MLD^{+}21]. Top- [MLD^{+}21]. Topic [HC16, YPL^{+}16]. Topics [RSvO^{+}22]. Touch [CNI15, EGQ16]. Touchscreen [ZGZ^{+}16]. Traceability [SF23]. tracking [CLRT20, MNL^{+}23]. Tractable [RNA^{+}23]. Traffic [Riv14]. Training [MBD^{+}15, PASR^{+}16, SLVV23, ZLG22]. Traits [KW16, SHKM19, TBK^{+}20, YC23].

Trajectories [CMGM18]. Transactions [KGCC18, JR11]. Transfer [MLD^{+}21, PYDM16]. Transferability [CSO^{+}22]. Transparency [CSO^{+}22, SW20]. Tree [GGC^{+}19, SKS^{+}24]. tree-emoji [SKS^{+}24]. Tree-in-the-loop [SKS^{+}24]. TribalGram [AYL^{+}22]. Tribe [AYL^{+}22]. Triggering [KR14]. Trust [AHJR18, BPUK18, CSO^{+}22, CWZF24, CLS^{+}18, JJT24, JCL^{+}19, KGCC18, LDM23, MJJT24, WRH18, WHLZ18]. Trust-Based [WHLZ18]. Trusting [ZMLY19]. Trustworthy [Shn20]. Truth [DAW18]. Tuberculosis [JIKB22]. turn [JFNY13]. turn-taking [JFNY13]. Tutorial [Gil15]. Tutoring [RHS19]. Twitter [OBG^{+}19]. Type [HBZ23].

Ubiquitously [MHC^{+}18]. UI [FJZ^{+}22]. Uncertainty [CHST23]. Understanding [AL23, BPUK18, BEH22, GZGW23, KSP^{+}20, RD21, SHA^{+}15b, SHA^{+}15a, ZFM^{+}24, ZTSH23]. Unified [JCL^{+}19]. Unify [BHS^{+}21]. Universal [SSL^{+}19]. Unobtrusive [APY^{+}20]. Unstructured [HMH^{+}19]. Unsupervised [YSD15]. Updatable [PCNB17]. Urban [LBC^{+}21, PSMO18]. Usability [FLT20, YPL^{+}16]. Use [BPUP22, HZKW21, RD21]. User [AL22, AL23, BEC^{+}20, BSKC21, BBQ20, BCKC20, BLO22, BEH22, CRO^{+}16, CMPS17, CHST23, CDF19, CYW19, COAR16, CLRT20, DSV17, DSEV19, DK18, EVB^{+}20, FB18, FWY^{+}18, JWK14, KMAM20, LPBG19, LBC^{+}21, NRB^{+}22, NCWC17, OPM17, DBTV20, SLMK18, SV22, SLH^{+}21, SCC14, WDN24, YC23, Yan21, YVA^{+}24, ZTD^{+}24, INN13, KSW^{+}11, LDM23]. User-Adaptive [SLH^{+}21, LPBG19]. User-Centered [FB18, Yan21]. User-Centred [SV22]. User-Controllable [DBTV20]. User-defined [BLO22, LBC^{+}21]. User-Driven [DSV17].

User-generated [DSEV19]. User-Reported [EVB^{+}20]. Users [BEH22, LDM23, SF23, SW20, VPB16, YNK^{+}22, NSA^{+}22]. Using [AYL^{+}22, AHJR18, APY^{+}20, BLV14, BDO^{+}16, CDK^{+}18, DPR^{+}22, FB18, HRAR18, IOKY16a, IOKY16b, KKS22, MNL^{+}23, PSMO18, PLBG19, SYS^{+}15, SG12, TBK^{+}20, YYI^{+}16, AFFB14, RNA^{+}23, SGS^{+}12]. USMART [YSD15]. Utility [MVH20]. Utilizing [PCK^{+}13].

VADAF [MWP^{+}21]. Valence [MKW21]. Valence-Arousal [MKW21]. Validating [ZTSH23]. Validation [ACMH23, CMGM18]. Value [RAMZ^{+}17]. Values [KMAM20]. Variation [CMTB15]. Varying [MKW21]. Vehicle [DM17]. Vehicles [HWM^{+}22]. VERB [RDP^{+}24]. Via [ITB19, CLJ21, FLJA23, GSG^{+}24, SV22, YVA^{+}24]. Video [AGH^{+}23, BDO^{+}16, RFR^{+}24, SCKOGP15, TBK^{+}20]. VideoMob [GRNW15]. Videos [RNA^{+}23, TTL^{+}22, YLVE12]. View [TRO^{+}19]. Violent [Hol18]. Virtual [LCKM23, PLBW17, RSPT^{+}23, TSSL15, WZB^{+}16, ZMLY19, GCDN11]. VisForum [FWY^{+}18]. Vision [BLV14, ITB19]. Visitor [DK15]. Visual [AGH^{+}23, CCM14, DSV17, DPS^{+}19, EGQ16, FLJA23, FWY^{+}18, GGC^{+}19, GSG^{+}24, HSS^{+}21, JLH^{+}23, KLC^{+}21, KCC^{+}19, KPSK17, LWFM23, LC18, NSMM18, SJS^{+}21, SMM20, SLMK18, SKKM15, SLH^{+}21, SKG^{+}18, TJLO20, TB20, TVA^{+}21, VPB16, WDN24, ZLW17, ZTSH23, AFFB14]. Visual-Spatial [ZLW17]. Visualization [AGH^{+}23, AG13, BLIC21, CLJ21, CLRT20, GSC^{+}17, LBC^{+}21, MWP^{+}21, Riv14, SLH^{+}21, SCC14]. Visualizations [BLIC21, LZW^{+}24, MVT16, TTL^{+}22]. Visualizer [BSKB21]. Visualizing [MHC^{+}18, RDP^{+}24, WSL^{+}18]. Visuals [ABJ15]. VizRec [MVT16]. Vlog [SCKOGP15]. Voice [BPUP22]. Volume [MSD^{+}16]. Vs [VPB16, VBCC20]. Vulnerability [LWFM23].

Want [SW20]. watching [YLVE12]. Wearable [MBD^{+}15, YNK^{+}22]. Wearables [HRAR18, ML19, PLBG19]. Web [AMN^{+}13, FLJA23, WBL^{+}13]. Well [MDR^{+}22]. Well-being [MDR^{+}22]. Wellbeing [KLC^{+}22]. Who [IOKY16a, IOKY16b]. Why-oriented [KSW^{+}11]. Wikipedia [DSEV19, SSL^{+}19]. Wikipedia-Based [SSL^{+}19]. Wild [MDR^{+}22]. Will [IOKY16a, IOKY16b]. within [BHNP20]. Wizard [BL15]. Word [KD20, RDP^{+}24]. Work [APY^{+}20, CWZF24]. Worker [KKS22]. Workflows [DFR15, JJL16]. Workload [CLS^{+}18]. Workspace [KKS22]. world [CAB^{+}13]. would [CWZF24]. WOZ [BL15]. Wristband [LPBG19]. Writing [TVCL16].

XAutoML [ZTSH23].

Young [KMO22, RD21].

References

Alaoui:2015:IVM

[ABJ15] Sarah Fdili Alaoui, Frederic Bevilacqua, and Christian Jacquemin. Interactive visuals as metaphors for dance movement qualities. ACM Transactions on Interactive Intelligent Systems (TIIS), 5 (3):13:1–13:??, October 2015. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Andre:2013:ISS

[AC13] Elisabeth André and Joyce Chai. Introduction to the special section on eye gaze and conversation. ACM Transactions on Interactive Intelligent Systems (TIIS), 3(2):10:1–10:??, July 2013. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Aguirre:2023:CTC

[ACMH23] Carlos Aguirre, Shiye Cao, Amama Mahmood, and Chien-Ming Huang. Crowdsourcing thumbnail captions: Data collection and validation. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(3): 14:1–14:??, September 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3589346.

Anderson:2020:MMM

[ADS^{+}20] Andrew Anderson, Jonathan Dodge, Amrita Sadarangani, Zoe Juozapaitis, Evan Newman, Jed Irvine, Souti Chattopadhyay, Matthew Olson, Alan Fern, and Margaret Burnett. Mental models of mere mortals with explanations of reinforcement learning. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(2): 15:1–15:37, June 2020. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

tronic). URL https://dl.

                                        acm.org/doi/abs/10.1145/
                                        3366485.

Alhejaili:2022:ELF

[AF22] Abdullah Alhejaili and Shaheen Fatima. Expressive latent feature modelling for explainable matrix factorisation-based recommender systems. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(3):20:1–20:30, September 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3530299.

Afergan:2018:DR

[Afe18]  Daniel Afergan.  Distin-

guished reviewers. ACM Transactions on Interactive Intelligent Systems (TIIS), 8(4):32:1–32:??, November 2018. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3283374.

Apostolopoulos:2014:IOL

[AFFB14] Ilias Apostolopoulos, Navid Fallah, Eelke Folmer, and Kostas E. Bekris. Integrated online localization and navigation for people with visual impairments using smart phones. ACM Transactions on Interactive Intelligent Systems (TIIS), 3(4): 21:1–21:??, January 2014. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Amir:2013:PRV

[AG13] Ofra Amir and Ya’akov (Kobi)

Gal. Plan recognition and visualization in exploratory learning environments. ACM Transactions on Interactive Intelligent Systems (TIIS), 3 (3):16:1–16:??, October 2013. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Afzal:2023:VVA

[AGH^{+}23] Shehzad Afzal, Sohaib Ghani, Mohamad Mazen Hittawe, Sheikh Faisal Rashid, Omar M. Knio, Markus Hadwiger, and Ibrahim Hoteit. Visualization and visual analytics approaches for image and video datasets: a survey. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(1):5:1–5:??, March 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3576935.

Akash:2018:CMS

[AHJR18] Kumar Akash, Wan-Lin Hu, Neera Jain, and Tahira Reid. A classification model for sensing human trust in machines using EEG and GSR. ACM Transactions on Interactive Intelligent Systems (TIIS), 8(4):27:1– 27:??, November 2018. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3132743.

Ahn:2020:PIP

[AL20] Yongsu Ahn and Yu-Ru Lin. PolicyFlow: Interpreting policy diffusion in context. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(2):13:1–13:23, June 2020. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3385729.

Ang:2022:LSR

[AL22] Gary Ang and Ee-Peng Lim. Learning semantically rich network-based multi-modal mobile user interface embeddings. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(4): 34:1–34:??, December 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3533856.

Ang:2023:LUU

[AL23] Gary Ang and Ee-Peng Lim. Learning and understanding user interface semantics from heterogeneous networks with multimodal and positional attributes. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(3):12:1–12:??, September 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3578522.

Amershi:2013:LAW

[AMN^{+}13] Saleema Amershi, Jalal Mahmud, Jeffrey Nichols, Tessa Lau, and German Attanasio Ruiz. LiveAction: Automating Web task model generation. ACM Transactions on Interactive Intelligent Systems (TIIS), 3 (3):14:1–14:??, October 2013. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Andrews:2012:SPP

[And12] Pierre Y. Andrews. System personality and persuasion in human-computer dialogue. ACM Transactions on Interactive Intelligent Systems (TIIS), 2(2):12:1–12:??, June 2012. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Avrahami:2020:UAR

[APY^{+}20] Daniel Avrahami, Mitesh Patel, Yusuke Yamaura, Sven Kratz, and Matthew Cooper. Unobtrusive activity recognition and position estimation for work surfaces using RFradar sensing. ACM Transactions on Interactive Intelligent Systems (TIIS), 10 (1):11:1–11:28, January 2020. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3241383.

Aslan:2016:DEM

[AUMT16] Ilhan Aslan, Andreas Uhl, Alexander Meschtscherjakov, and Manfred Tscheligi. Design and exploration of mid-air authentication gestures. ACM Transactions on Interactive Intelligent Systems (TIIS), 6(3):23:1–23:??, October 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Ahn:2022:TCI

[AYL^{+}22] Yongsu Ahn, Muheng Yan, Yu-Ru Lin, Wen-Ting Chung, and Rebecca Hwa. Tribe or not? Critical inspection of group differences using TribalGram. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(1): 5:1–5:34, March 2022. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3484509.

Billinghurst:2020:SIH

[BBQ20] Mark Billinghurst, Margaret Burnett, and Aaron Quigley. Special issue on highlights of ACM Intelligent User Interface (IUI) 2018. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(1):1:1–1:3, January 2020.

CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3357206.

Bulling:2015:ISI

[BBT^{+}15] Andreas Bulling, Ulf Blanke, Desney Tan, Jun Rekimoto, and Gregory Abowd. Introduction to the special issue on activity recognition for interaction. ACM Transactions on Interactive Intelligent Systems (TIIS), 4 (4):16:1–16:??, January 2015. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Brdiczka:2020:ISI

[BCKC20] Oliver Brdiczka, Duen Horng Chau, Minsuk Kahng, and Gaëlle Calvary. Introduction to the special issue on highlights of ACM Intelligent User Interface (IUI) 2019. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(4):27:1–27:2, December 2020. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL

https://dl.acm.org/doi/
10.1145/3429946.

Bosch:2016:UVA

[BDO^{+}16] Nigel Bosch, Sidney K.

D’mello, Jaclyn Ocumpaugh, Ryan S. Baker, and Valerie Shute. Using video to automatically detect learner affect in computer-enabled classrooms. ACM Transactions on Interactive Intelligent Systems (TIIS), 6 (2):17:1–17:??, August 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Bagheri:2020:ACE

[BEC^{+}20] Elahe Bagheri, Pablo G. Esteban, Hoang-Long Cao, Albert De Beir, Dirk Lefeber, and Bram Vanderborght. An autonomous cognitive empathy model responsive to users’ facial emotion expressions. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(3): 20:1–20:23, November 2020. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3341198.

Buschek:2022:HSU

[BEH22] Daniel Buschek, Malin Eiband, and Heinrich Hussmann. How to support users in understanding intelligent systems? an analysis and conceptual framework of user questions considering user mindsets, involvement, and knowledge outcomes. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(4):29:1–29:??, December 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3519264.

Berkovsky:2012:IIF

[BFOK12] Shlomo Berkovsky, Jill Freyne, and Harri Oinas-Kukkonen. Influencing individually: Fusing personalization and persuasion. ACM Transactions on Interactive Intelligent Systems (TIIS), 2(2): 9:1–9:??, June 2012. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Banisetty:2021:SAN

[BFY^{+}21] Santosh Balajee Banisetty, Scott Forer, Logan Yliniemi, Monica Nicolescu, and David Feil-Seifer. Socially aware navigation: a nonlinear multi-objective optimization approach. ACM Transactions on Interactive Intelligent Systems (TIIS), 11(2):15:1–15:26, July 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3453445.

Broden:2020:BBE

[BHNP20] Björn Brodén, Mikael Hammar, Bengt J. Nilsson, and Dimitris Paraschakis. A bandit-based ensemble framework for exploration/ exploitation of diverse recommendation components: an experimental study within ecommerce. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(1): 4:1–4:32, January 2020. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3237187.

Bernard:2021:TPM

[BHS^{+}21] Jürgen Bernard, Marco Hutter, Michael Sedlmair, Matthias Zeppelzauer, and Tamara Munzner. A taxonomy of property measures to unify active learning and human-centered approaches to data labeling. ACM Transactions on Interactive Intelligent Systems (TIIS), 11 (3–4):20:1–20:42, December 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3439333.

Bott:2015:WRW

[BL15] Jared N. Bott and Joseph J. Laviola Jr. The WOZ Recognizer: a Wizard of Oz sketch recognition system. ACM Transactions on Interactive Intelligent Systems (TIIS), 5 (3):15:1–15:??, October 2015. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Barral:2021:EAG

[BLIC21] Oswald Barral, Sébastien Lallé, Alireza Iranpour, and Cristina Conati. Effect of adaptive guidance and visualization literacy on gaze attentive behaviors and sequential patterns on magazinestyle narrative visualizations.

ACM Transactions on Interactive Intelligent Systems (TIIS), 11(3–4):28:1–28:46, December 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL

https://dl.acm.org/doi/
10.1145/3447992.

Bruckner:2022:LGC

[BLO22] Lukas Brückner, Luis A.

Leiva, and Antti Oulasvirta. Learning GUI completions with user-defined constraints. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(1):6:1–6:40, March 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL

https://dl.acm.org/doi/
10.1145/3490034.

Benotti:2014:INL

[BLV14] Luciana Benotti, Tessa Lau, and Martín Villalba. Interpreting natural language instructions using language, vision, and behavior. ACM Transactions on Interactive Intelligent Systems (TIIS), 4 (3):13:1–13:??, August 2014. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Baur:2015:CAA

[BMD^{+}15] Tobias Baur, Gregor Mehlmann, Ionut Damian, Florian Lingenfelser, Johannes Wagner, Birgit Lugrin, Elisabeth André, and Patrick Gebhard. Context-aware automated analysis and annotation of social human–agent interactions. ACM Transactions on Interactive Intelligent Systems (TIIS), 5(2): 11:1–11:??, July 2015. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Baker:2018:TUT

[BPUK18] Anthony L. Baker, Elizabeth K. Phillips, Daniel Ullman, and Joseph R. Keebler. Toward an understanding of trust repair in human-robot interaction: Current research and future directions. ACM Transactions on Interactive Intelligent Systems (TIIS), 8(4):30:1–30:??, November 2018. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3181671.

Brewer:2022:ESO

[BPUP22] Robin Brewer, Casey Pierce, Pooja Upadhyay, and Leeseul Park. An empirical study of older adult’s voice assistant use for health information seeking. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(2): 13:1–13:32, June 2022. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3484507.

Beck:2012:EBL

[BSBC12] Aryel Beck, Brett Stevens, Kim A. Bard, and Lola Cañamero. Emotional body language displayed by artificial agents. ACM Transactions on Interactive Intelligent Systems (TIIS), 2(1): 2:1–2:??, March 2012. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Bhattacharya:2021:RTI

[BSKB21] Samit Bhattacharya, Viral Bharat Shah, Krishna Kumar, and Ujjwal Biswas. A real-time interactive visualizer for large classroom. ACM Transactions on Interactive Intelligent Systems (TIIS), 11(1):6:1–6:26, April 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3418529.

Bessghaier:2021:DSA

[BSKC21] Narjes Bessghaier, Makram Soui, Christophe Kolski, and Mabrouka Chouchane. On the detection of structural aesthetic defects of Android mobile user interfaces with a metrics-based tool. ACM Transactions on Interactive Intelligent Systems (TIIS), 11(1):3:1–3:27, April 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3410468.

Console:2013:ISN

[CAB^{+}13] Luca Console, Fabrizio Antonelli, Giulia Biamino, Francesca Carmagnola, Federica Cena, Elisa Chiabrando, Vincenzo Cuciti, Matteo Demichelis, Franco Fassio, Fabrizio Franceschi, Roberto Furnari, Cristina Gena, Marina Geymonat, Piercarlo Grimaldi, Pierluige Grillo, Silvia Likavec, Ilaria Lombardi, Dario Mana, Alessandro Marcengo, Michele Mioli, Mario Mirabelli, Monica Perrero, Claudia Picardi, Federica Protti, Amon Rapp, Rossana Simeoni, Daniele Theseider Dupré, Ilaria Torre, Andrea Toso, Fabio Torta, and Fabiana Vernero. Interacting with social networks of intelligent things and people in the world of gastronomy. ACM Transactions on Interactive Intelligent Systems (TIIS), 3 (1):4:1–4:??, April 2013. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Chan:2014:RCT

[CCM14] Yu-Hsuan Chan, Carlos D. Correa, and Kwan-Liu Ma. Regression Cube: a technique for multidimensional visual exploration and interactive pattern finding. ACM Transactions on Interactive Intelligent Systems (TIIS), 4 (1):7:1–7:??, April 2014. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Chen:2019:SIH

[CDF19] Fang Chen, Carlos Duarte, and Wai-Tat Fu. Special issue on highlights of ACM Intelligent User Interface (IUI) 2017. ACM Transactions on Interactive Intelligent Systems (TIIS), 9(2–3): 7:1–7:??, April 2019. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3301292.

Chen:2013:HDM

[CdGF^{+}13] Li Chen, Marco de Gemmis, Alexander Felfernig, Pasquale Lops, Francesco Ricci, and Giovanni Semeraro. Human decision making and recommender systems. ACM Transactions on Interactive Intelligent Systems (TIIS), 3(3):17:1–17:??, October 2013. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Chen:2018:UML

[CDK^{+}18] Nan-Chen Chen, Margaret Drouhard, Rafal Kocielnik, Jina Suh, and Cecilia R. Aragon. Using machine learning to support qualitative coding in social science: Shifting the focus to ambiguity. ACM Transactions on Interactive Intelligent Systems (TIIS), 8(2): 9:1–9:??, July 2018. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Cuayahuitl:2014:ISI

[CFD^{+}14] Heriberto Cuayáhuitl, Lutz Frommberger, Nina Dethlefs, Antoine Raux, Mathew Marge, and Hendrik Zender. Introduction to the special issue on machine learning for multiple modalities in interactive systems and robots. ACM Transactions on Interactive Intelligent Systems (TIIS), 4(3): 12e:1–12e:??, October 2014. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Cena:2017:ISI

[CGHS17] F. Cena, C. Gena, G. J. Houben, and M. Strohmaier. Introduction to the Special Issue on Big Personal Data in Interactive Intelligent Systems. ACM Transactions on Interactive Intelligent Systems (TIIS), 7(2): 6:1–6:??, July 2017. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Cremonesi:2012:IPP

[CGT12] Paolo Cremonesi, Franca Garzotto, and Roberto Turrin. Investigating the persuasion potential of recommender systems from a quality perspective: an empirical study. ACM Transactions on Interactive Intelligent Systems (TIIS), 2(2): 11:1–11:??, June 2012. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Cau:2023:EAL

[CHST23] Federico Maria Cau, Hanna Hauptmann, Lucio Davide Spano, and Nava Tintarev. Effects of AI and logic-style explanations on users’ decisions under different levels of uncertainty. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(4):22:1–22:??, December 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3588320.

Cuayahuitl:2014:NHR

[CKKD14] Heriberto Cuayáhuitl, Ivana Kruijff-Korbayová, and Nina Dethlefs. Nonstrict hierarchical reinforcement learning for interactive systems and robots. ACM Transactions on Interactive Intelligent Systems (TIIS), 4 (3):15:1–15:??, October 2014. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Chen:2021:PPR

[CLJ21] Xiaoyu Chen, Nathan Lau, and Ran Jin. PRIME: a personalized recommender system for information visualization methods via extended matrix completion. ACM Transactions on Interactive Intelligent Systems (TIIS), 11(1):7:1–7:30, April 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3366484.

Castellano:2014:CSA

[CLP^{+}14] Ginevra Castellano, Iolanda Leite, André Pereira, Carlos Martinho, Ana Paiva, and Peter W. Mcowan. Contextsensitive affect recognition for a robotic game companion. ACM Transactions on Interactive Intelligent Systems (TIIS), 4(2): 10:1–10:??, July 2014. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Conati:2020:CCI

[CLRT20] Cristina Conati, Sébastien Lallé, Md Abed Rahman, and Dereck Toker. Comparing and combining interaction data and eye-tracking data for the real-time prediction of user cognitive abilities in visualization tasks. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(2):12:1–12:41, June 2020. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3301400.

Chien:2018:ECT

[CLS^{+}18] Shih-Yi Chien, Michael Lewis, Katia Sycara, Jyi-Shane Liu, and Asiye Kumru. The effect of culture on trust in automation: Reliability and workload. ACM Transactions on Interactive Intelligent Systems (TIIS), 8(4): 29:1–29:??, November 2018. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3230736.

Carreno-Medrano:2018:PVG

[CMGM18] Pamela Carreno-Medrano, Sylvie Gibet, and Pierre-FranÇois Marteau. Perceptual validation for the generation of expressive movements from end-effector trajectories. ACM Transactions on Interactive Intelligent Systems (TIIS), 8(3): 24:1–24:??, August 2018. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Cassavia:2017:DUB

[CMPS17] Nunziato Cassavia, Elio Masciari, Chiara Pulice, and Domenico Saccà. Discovering user behavioral features to enhance information search on big data. ACM Transactions on Interactive Intelligent Systems (TIIS), 7(2): 7:1–7:??, July 2017. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Caramiaux:2015:AGR

[CMTB15] Baptiste Caramiaux, Nicola Montecchio, Atau Tanaka, and Frédéric Bevilacqua. Adaptive gesture recognition with variation estimation for interactive systems. ACM Transactions on Interactive Intelligent Systems (TIIS), 4 (4):18:1–18:??, January 2015. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Cooney:2015:AIS

[CNI15] Martin Cooney, Shuichi Nishio, and Hiroshi Ishiguro. Affectionate interaction with a small humanoid robot capable of recognizing social touch behavior. ACM Transactions on Interactive Intelligent Systems (TIIS), 4 (4):19:1–19:??, January 2015. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Clark:2016:MAA

[COAR16] Leigh Clark, Abdulmalik Ofemile, Svenja Adolphs, and Tom Rodden. A multimodal approach to assessing user experiences with agent helpers. ACM Transactions on Interactive Intelligent Systems (TIIS), 6(4): 29:1–29:??, December 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Chen:2014:EPM

[CQW^{+}14] Yingjie Victor Chen, Zhenyu Cheryl Qian, Robert Woodbury, John Dill, and Chris D. Shaw. Employing a parametric model for analytic provenance. ACM Transactions on Interactive Intelligent Systems (TIIS), 4(1): 6:1–6:??, April 2014. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Chen:2012:MBI

[CRC^{+}12] Fang Chen, Natalie Ruiz, Eric Choi, Julien Epps, M. Asif Khawaja, Ronnie Taib, Bo Yin, and Yang Wang. Multimodal behavior and interaction as indicators of cognitive load. ACM Transactions on Interactive Intelligent Systems (TIIS), 2(4):22:1–22:??, December 2012. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Cafaro:2016:EIA

[CRO^{+}16] Angelo Cafaro, Brian Ravenet, Magalie Ochs, Hannes Högni Vilhjálmsson, and Catherine Pelachaud. The effects of interpersonal attitude of a group of agents on user’s presence and proxemics behavior. ACM Transactions on Interactive Intelligent Systems (TIIS), 6 (2):12:1–12:??, August 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Castellano:2012:ISI

[CRP^{+}12] Ginevra Castellano, Laurel D. Riek, Christopher Peters, Kostas Karpouzis, Jean-Claude Martin, and Louis-Philippe Morency. Introduction to the special issue on affective interaction in natural environments. ACM Transactions on Interactive Intelligent Systems (TIIS), 2(1): 1:1–1:??, March 2012. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Callaway:2014:EMD

[CSD14] Charles Callaway, Oliviero Stock, and Elyon Dekoven. Experiments with mobile drama in an instrumented museum for inducing conversation in small groups. ACM Transactions on Interactive Intelligent Systems (TIIS), 4 (1):2:1–2:??, April 2014. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Carberry:2012:AMA

[CSM^{+}12] Sandra Carberry, Stephanie Elzer Schwartz, Kathleen Mccoy, Seniz Demir, Peng Wu, Charles Greenbacker, Daniel Chester, Edward Schwartz, David Oliver, and Priscilla Moraes. Access to multimodal articles for individuals with sight impairments. ACM Transactions on Interactive Intelligent Systems (TIIS), 2(4):21:1–21:??, December 2012. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Caldwell:2022:ANR

[CSO^{+}22] Sabrina Caldwell, Penny Sweetser, Nicholas O’Donnell, Matthew J. Knight, Matthew Aitchison, Tom Gedeon, Daniel Johnson, Margot Brereton, Marcus Gallagher, and David Conroy. An agile new research framework for hybrid human-AI teaming: Trust, transparency, and transferability. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(3):17:1–17:36, September 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3514257.

Cheng:2024:IWW

[CWZF24] Ruijia Cheng, Ruotong Wang, Thomas Zimmermann, and Denae Ford. “It would work for me too”: How online communities shape software developers’ trust in AI-powered code generation tools. ACM Transactions on Interactive Intelligent Systems (TIIS), 14(2): 11:1–11:??, June 2024. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3651990.

Chen:2019:UES

[CYW19] Li Chen, Dongning Yan, and Feng Wang. User evaluations on sentimentbased recommendation explanations. ACM Transactions on Interactive Intelligent Systems (TIIS), 9(4): 20:1–20:??, December 2019. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3282878.

Dumitrache:2018:CGT

[DAW18] Anca Dumitrache, Lora Aroyo, and Chris Welty. Crowdsourcing ground truth for medical relation extraction. ACM Transactions on Interactive Intelligent Systems (TIIS), 8(2):11:1– 11:??, July 2018. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Sciascio:2020:RUC

[DBTV20] Cecilia Di Sciascio, Peter Brusilovsky, Christoph Trattner, and Eduardo Veas. A roadmap to user-controllable social exploratory search. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(1):8:1–8:38, January 2020. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL

                                        https://dl.acm.org/doi/
                                        abs/10.1145/3241382.

Dominguez:2020:AHA

[DDGMP20] Vicente Dominguez, Ivania Donoso-Guzmán, Pablo Messina, and Denis Parra. Algorithmic and HCI aspects for explaining recommendations of artistic images. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(4):30:1–30:31, December 2020. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3369396.

DeCarolis:2015:ILD

[DFR15] Berardina De Carolis, Stefano Ferilli, and Domenico Redavid. Incremental learning of daily routines as workflows in a Smart home environment. ACM Transactions on Interactive Intelligent Systems (TIIS), 4 (4):20:1–20:??, January 2015. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Dmello:2012:AAA

[DG12] Sidney D’mello and Art Graesser. AutoTutor and Affective Autotutor: Learning by talking with cognitively and emotionally intelligent computers that talk back. ACM Transactions on Interactive Intelligent Systems (TIIS), 2(4):23:1–23:??, December 2012. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Dardard:2016:ACL

[DGG16] Floriane Dardard, Giorgio Gnecco, and Donald Glowinski. Automatic classification of leading interactions in a string quartet. ACM Transactions on Interactive Intelligent Systems (TIIS), 6(1): 5:1–5:??, May 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Davis:2017:QCC

[DHS^{+}17] N. Davis, C. Hsiao, K. Y. Singh, B. Lin, and B. Magerko. Quantifying collaboration with a co-creative drawing agent. ACM Transactions on Interactive Intelligent Systems (TIIS), 7(4):19:1–19:??, December 2017. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Dinakar:2012:CSR

[DJH^{+}12] Karthik Dinakar, Birago Jones, Catherine Havasi, Henry Lieberman, and Rosalind Picard. Common sense reasoning for detection, prevention, and mitigation of cyberbullying. ACM Transactions on Interactive Intelligent Systems (TIIS), 2(3): 18:1–18:??, September 2012. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Dim:2015:ADS

[DK15] Eyal Dim and Tsvi Kuflik. Automatic detection of social behavior of museum visitor pairs. ACM Transactions on Interactive Intelligent Systems (TIIS), 4 (4):17:1–17:??, January 2015. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Dudley:2018:RUI

[DK18] John J. Dudley and Per Ola Kristensson. A review of user interface design for interactive machine learning. ACM Transactions on Interactive Intelligent Systems (TIIS), 8 (2):8:1–8:??, July 2018. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Dodge:2021:AAR

[DKI^{+}21] Jonathan Dodge, Roli Khanna, Jed Irvine, Kin ho Lam, Theresa Mai, Zhengxian Lin, Nicholas Kiddle, Evan Newman, Andrew Anderson, Sai Raja, Caleb Matthews, Christopher Perdriau, Margaret Burnett, and Alan Fern. After-action review for AI (AAR/AI). ACM Transactions on Interactive Intelligent Systems (TIIS), 11 (3–4):29:1–29:35, December 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3453173.

Deng:2015:ESA

[DLMC15] James J. Deng, Clement H. C. Leung, Alfredo Milani, and Li Chen. Emotional states associated with music: Classification, prediction of changes, and consideration in recommendation. ACM Transactions on Interactive Intelligent Systems (TIIS), 5(1):4:1–4:??, March 2015. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Duncan:2017:ESC

[DM17] Brittany A. Duncan and Robin R. Murphy. Effects of speed, cyclicity, and dimensionality on distancing, time, and preference in human–aerial vehicle interactions. ACM Transactions on Interactive Intelligent Systems (TIIS), 7(3): 13:1–13:??, October 2017. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Douer:2021:TMS

[DM21] Nir Douer and Joachim Meyer. Theoretical, measured, and subjective responsibility in aided decision making. ACM Transactions on Interactive Intelligent Systems (TIIS), 11(1): 5:1–5:37, April 2021. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3425732.

Dodson:2013:ELA

[DMGG13] Thomas Dodson, Nicholas Mattei, Joshua T. Guerin, and Judy Goldsmith. An English-language argumentation interface for explanation generation with Markov decision processes in the domain of academic advising. ACM Transactions on Interactive Intelligent Systems (TIIS), 3 (3):18:1–18:??, October 2013. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Das:2023:EAR

[DNV^{+}23] Devleena Das, Yasutaka Nishimura, Rajan P. Vivek, Naoto Takeda, Sean T. Fish, Thomas Plötz, and Sonia Chernova. Explainable activity recognition for smart home systems. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(2): 7:1–7:??, June 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL

https://dl.acm.org/doi/
10.1145/3561533.

Dotti:2020:BCA

[DPA20] Dario Dotti, Mirela Popa, and Stylianos Asteriadis. Being the center of attention: a person-context CNN framework for personality recognition. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(3): 19:1–19:20, November 2020. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3338245.

Das:2022:DRD

[DPR^{+}22] Kapotaksha Das, Michalis Papakostas, Kais Riani, Andrew Gasiorowski, Mohamed Abouelenien, Mihai Burzo, and Rada Mihalcea. Detection and recognition of driver distraction using multimodal signals. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(4): 33:1–33:??, December 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3519267.

Du:2019:EVA

[DPS^{+}19] Fan Du, Catherine Plaisant, Neil Spring, Kenyon Crowley, and Ben Shneiderman. EventAction: a visual analytics approach to explainable recommendation for event sequences. ACM Transactions on Interactive Intelligent Systems (TIIS), 9(4): 21:1–21:??, December 2019. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3301402.

DiSciascio:2019:IQA

[DSEV19] Cecilia Di Sciascio, David Strohmaier, Marcelo Errecalde, and Eduardo Veas. Interactive quality analytics of user-generated content: an integrated toolkit for the case of Wikipedia. ACM Transactions on Interactive Intelligent Systems (TIIS), 9(2–3): 13:1–13:??, April 2019. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3150973.

DiSciascio:2017:SES

[DSV17] Cecilia Di Sciascio, Vedran Sabol, and Eduardo Veas. Supporting exploratory search with a visual userdriven approach. ACM Transactions on Interactive Intelligent Systems (TIIS), 7(4):18:1–18:??, December 2017. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Eagle:2022:DKW

[EBB^{+}22] Tessa Eagle, Conrad Blau, Sophie Bales, Noopur Desai, Victor Li, and Steve Whittaker. “I don’t know what you mean by ‘I am anxious”’: a new method for evaluating conversational agent responses to standardized mental health inputs for anxiety and depression. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(2):12:1–12:23, June 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3488057.

Emamgholizadeh:2024:PGC

[EDR24] Hanif Emamgholizadeh, Amra Delić, and Francesco Ricci. Predicting group choices from group profiles. ACM Transactions on Interactive Intelligent Systems (TIIS), 14(1):7:1–7:??, March 2024. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3639710.

El-Glaly:2016:RWY

[EGQ16] Yasmine N. El-Glaly and Francis Quek. Read what you touch with intelligent audio system for non-visual interaction. ACM Transactions on Interactive Intelligent Systems (TIIS), 6 (3):24:1–24:??, October 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Eiband:2020:MAE

[EVB^{+}20] Malin Eiband, Sarah Theres Völkel, Daniel Buschek, Sophia Cook, and Heinrich Hussmann. A method and analysis to elicit userreported problems in intelligent everyday applications. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(4):28:1–28:27, December 2020. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3370927.

Eyben:2012:MAC

[EWS12] Florian Eyben, Martin Wöllmer, and Björn Schuller. A multitask approach to continuous five-dimensional affect sensing in natural speech. ACM Transactions on Interactive Intelligent Systems (TIIS), 2(1):6:1–6:??, March 2012. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Francoise:2018:MSM

[FB18] Jules Françoise and Frédéric Bevilacqua. Motion-sound mapping through interaction: an approach to user-centered design of auditory feedback using machine learning. ACM Transactions on Interactive Intelligent Systems (TIIS), 8(2):16:1–16:??, July 2018. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Fang:2014:CLM

[FBC14] Yi Fang, Ziad Al Bawab, and Jean-François Crespo. Collaborative language models for localized query prediction. ACM Transactions on Interactive Intelligent Systems (TIIS), 4(2): 9:1–9:??, July 2014. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Freyne:2013:RBP

[FBS13] Jill Freyne, Shlomo Berkovsky, and Gregory Smith. Rating bias and preference acquisition. ACM Transactions on Interactive Intelligent Systems (TIIS), 3(3): 19:1–19:??, October 2013. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Fiebrink:2018:ISI

[FG18] Rebecca Fiebrink and Marco Gillies. Introduction to the special issue on humancentered machine learning. ACM Transactions on Interactive Intelligent Systems (TIIS), 8(2):7:1–7:??, July 2018. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Feng:2022:AIA

[FJZ^{+}22] Sidong Feng, Minmin Jiang, Tingting Zhou, Yankun Zhen, and Chunyang Chen. Auto-Icon+: an automated end-to-end code generation tool for icon designs in UI development. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(4): 36:1–36:??, December 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3531065.

Ferdous:2023:EEW

[FLJA23] Javedul Ferdous, Hae-Na Lee, Sampath Jayarathna, and Vikas Ashok. Enabling efficient Web data-record interaction for people with visual impairments via proxy interfaces. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(3): 13:1–13:??, September 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3579364.

Fan:2020:ADU

[FLT20] Mingming Fan, Yue Li, and Khai N. Truong. Automatic detection of usability problem encounters in thinkaloud sessions. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(2): 16:1–16:24, June 2020. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3385732.

Fu:2018:VVA

[FWY^{+}18] Siwei Fu, Yong Wang, Yi Yang, Qingqing Bi, Fangzhou Guo, and Huamin Qu. VisForum: a visual analysis system for exploring user groups in online forums. ACM Transactions on Interactive Intelligent Systems (TIIS), 8(1):3:1–3:??, March 2018. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Gibet:2011:SSD

[GCDN11] Sylvie Gibet, Nicolas Courty, Kyle Duarte, and Thibaut Le Naour. The SignCom system for data-driven animation of interactive virtual signers: Methodology and evaluation. ACM Transactions on Interactive Intelligent Systems (TIIS), 1(1): 6:1–6:??, October 2011. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Guo:2019:VEA

[GGC^{+}19] Fangzhou Guo, Tianlong Gu, Wei Chen, Feiran Wu, Qi Wang, Lei Shi, and Huamin Qu. Visual exploration of air quality data with a time-correlation-partitioning tree based on information theory. ACM Transactions on Interactive Intelligent Systems (TIIS), 9 (1):4:1–4:??, February 2019. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3182187.

Gil:2021:AIM

[GGK^{+}21] Yolanda Gil, Daniel Garijo, Deborah Khider, Craig A. Knoblock, Varun Ratnakar, Maximiliano Osorio, Hernán Vargas, Minh Pham, Jay Pujara, Basel Shbita, Binh Vu, Yao-Yi Chiang, Dan Feldman, Yijun Lin, Hayley Song, Vipin Kumar, Ankush Khandelwal, Michael Steinbach, Kshitij Tayal, Shaoming Xu, Suzanne A. Pierce, Lissa Pearson, Daniel Hardesty-Lewis, Ewa Deelman, Rafael Ferreira Da Silva, Rajiv Mayani, Armen R. Kemanian, Yuning Shi, Lorne Leonard, Scott Peckham, Maria Stoica, Kelly Cobourn, Zeya Zhang, Christopher Duffy, and Lele Shu. Artificial intelligence for modeling complex systems: Taming the complexity of expert models to improve decision making. ACM Transactions on Interactive Intelligent Systems (TIIS), 11(2):11:1–11:49, July 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3453172.

Gil:2015:HTI

[Gil15] Yolanda Gil. Human tutorial instruction in the raw. ACM Transactions on Interactive Intelligent Systems (TIIS), 5(1):2:1–2:??, March 2015. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Gianni:2015:SRF

[GKP15] Mario Gianni, Geert-Jan M.

Kruijff, and Fiora Pirri. A stimulus-response framework for robot control. ACM Transactions on Interactive Intelligent Systems (TIIS), 4 (4):21:1–21:??, January 2015. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Giunchiglia:2013:ISS

[GR13] Fausto Giunchiglia and David Robertson. Introduction to the special section on Internet-scale human problem solving. ACM Transactions on Interactive Intelligent Systems (TIIS), 3(1): 1:1–1:??, April 2013. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Gil:2012:CCK

[GRC^{+}12] Yolanda Gil, Varun Ratnakar, Timothy Chklovski, Paul Groth, and Denny Vrandecic. Capturing common knowledge about tasks: Intelligent assistance for todo lists. ACM Transactions on Interactive Intelligent Systems (TIIS), 2(3): 15:1–15:??, September 2012. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Grenader:2015:VIA

[GRNW15] Emily Grenader, Danilo Gasques Rodrigues, Fernando Nos, and Nadir Weibel. The VideoMob interactive art installation connecting strangers through inclusive digital crowds. ACM Transactions on Interactive Intelligent Systems (TIIS), 5(2): 7:1–7:??, July 2015. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Gotz:2017:ACM

[GSC^{+}17] David Gotz, Shun Sun, Nan Cao, Rita Kundu, and Anne-Marie Meyer. Adaptive contextualization methods for combating selection bias during high-dimensional visualization. ACM Transactions on Interactive Intelligent Systems (TIIS), 7(4): 17:1–17:??, December 2017. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Guo:2024:TAN

[GSG^{+}24] Yi Guo, Danqing Shi, Mingjuan Guo, Yanqiu Wu, Nan Cao, and Qing Chen. Talk2Data: a natural language interface for exploratory visual analysis via question decomposition. ACM Transactions on Interactive Intelligent Systems (TIIS), 14(2):8:1–8:??, June 2024. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3643894.

Guo:2023:GGF

[GZGW23] Mengtian Guo, Zhilan Zhou, David Gotz, and Yue Wang. GRAFS: Graphical faceted search system to support conceptual understanding in exploratory search. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(2):9:1–9:??, June 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3588319.

He:2020:DDA

[HBKM20] Yangyang He, Paritosh Bahirat, Bart P. Knijnenburg, and Abhilash Menon. A data-driven approach to designing for privacy in household IoT. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(1): 10:1–10:47, January 2020. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3241378.

Hoey:2012:PSD

[HBP^{+}12] Jesse Hoey, Craig Boutilier, Pascal Poupart, Patrick Olivier, Andrew Monk, and Alex Mihailidis. People, sensors, decisions: Customizable and adaptive technologies for assistance in healthcare. ACM Transactions on Interactive Intelligent Systems (TIIS), 2(4):20:1–20:??, December 2012. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Hernandez-Bocanegra:2023:ERT

[HBZ23] Diana C. Hernandez-Bocanegra and Jürgen Ziegler. Explaining recommendations through conversations: Dialog model and the effects of interface type and degree of interactivity. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(2): 6:1–6:??, June 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL

https://dl.acm.org/doi/
10.1145/3579541.

Hoque:2016:ITM

[HC16] Enamul Hoque and Giuseppe Carenini. Interactive topic modeling for exploring asynchronous online conversations: Design and evaluation of ConVisIT. ACM Transactions on Interactive Intelligent Systems (TIIS), 6(1): 7:1–7:??, May 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Hsu:2020:SPE

[HCD^{+}20] Yen-Chia Hsu, Jennifer Cross, Paul Dille, Michael Tasota, Beatrice Dias, Randy Sargent, Ting-Hao (Kenneth) Huang, and Illah Nourbakhsh. Smell Pittsburgh: Engaging community citizen science for air quality. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(4):32:1–32:49, December 2020. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3369397.

Hiolle:2012:ECB

[HCDRB12] Antoine Hiolle, Lola Cañamero, Marina Davila-Ross, and Kim A. Bard. Eliciting caregiving behavior in dyadic human-robot attachment-like interactions. ACM Transactions on Interactive Intelligent Systems (TIIS), 2(1): 3:1–3:??, March 2012. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Hailpern:2020:HIP

[HHC20] Joshua Hailpern, Mark Huber, and Ronald Calvo. How impactful is presentation in email? The effect of avatars and signatures. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(3): 24:1–24:26, November 2020. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3345641.

Hoi:2011:AMK

[HJ11] Steven C. H. Hoi and Rong Jin. Active multiple kernel learning for interactive 3D object retrieval systems. ACM Transactions on Interactive Intelligent Systems (TIIS), 1(1):3:1–3:??, October 2011. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Harper:2016:MDH

[HK16] F. Maxwell Harper and Joseph A. Konstan. The MovieLens datasets: History and context. ACM Transactions on Interactive Intelligent Systems (TIIS), 5 (4):19:1–19:??, January 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Hammond:2022:SIH

[HKOT22] Tracy Hammond, Bart Knijnenburg, John O’Donovan, and Paul Taele. Special issue on highlights of IUI 2021: Introduction. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(4):25:1–25:??, December 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3561516.

Hammond:2018:JAA

[HKR^{+}18] Tracy Hammond, Shalini Priya Ashok Kumar, Matthew Runyon, Josh Cherian, Blake Williford, Swarna Keshavabhotla, Stephanie Valentine, Wayne Li, and Julie Linsey. It’s not just about accuracy: Metrics that matter when modeling expert sketching ability. ACM Transactions on Interactive Intelligent Systems (TIIS), 8 (3):19:1–19:??, August 2018. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Hezarjaribi:2019:HLL

[HMH^{+}19] Niloofar Hezarjaribi, Sepideh Mazrouee, Saied Hemati, Naomi S. Chaytor, Martine Perrigue, and Hassan Ghasemzadeh. Human-in-the-loop learning for personalized diet monitoring from unstructured mobile data. ACM Transactions on Interactive Intelligent Systems (TIIS), 9(4):23:1– 23:??, December 2019. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3319370.

Handler:2022:CIQ

[HMO22] Abram Handler, Narges Mahyar, and Brendan O’Connor. ClioQuery: Interactive queryoriented text analytics for comprehensive investigation of historical news archives. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(3):22:1– 22:49, September 2022. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3524025.

Holbrook:2018:CVI

[Hol18] Colin Holbrook. Cues of violent intergroup conflict diminish perceptions of robotic personhood. ACM Transactions on Interactive Intelligent Systems (TIIS), 8(4): 28:1–28:??, November 2018. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

tronic). URL https://dl.

acm.org/ft_gateway.cfm?
id=3181674.

Hammond:2011:RSM

[HP11] Tracy Hammond and Brandon Paulson. Recognizing sketched multistroke primitives. ACM Transactions on Interactive Intelligent Systems (TIIS), 1(1):4:1–4:??, October 2011. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Hossain:2018:ASM

[HRAR18] H. M. Sajjad Hossain, Sreenivasan R. Ramamurthy, Md Abdullah Al Hafiz Khan, and Nirmalya Roy. An active sleep monitoring framework using wearables. ACM Transactions on Interactive Intelligent Systems (TIIS), 8 (3):22:1–22:??, August 2018. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Hinterreiter:2021:PPE

[HSS^{+}21] Andreas Hinterreiter, Christian Steinparz, Moritz SchÖfl, Holger Stitz, and Marc Streit. Projection path explorer: Exploring visual patterns in projected decisionmaking paths. ACM Transactions on Interactive Intelligent Systems (TIIS), 11 (3–4):22:1–22:29, December 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3387165.

Higuchi:2020:LCD

[HTOB^{+}20] Keita Higuchi, Hiroki Tsuchida, Eshed Ohn-Bar, Yoichi Sato, and Kris Kitani. Learning context-dependent personal preferences for adaptive recommendation. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(3):23:1–23:26, November 2020. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3359755.

Hsieh:2022:ADA

[HWM^{+}22] Sheng-Jen Hsieh, Andy R. Wang, Anna Madison, Chad Tossell, and Ewart de Visser. Adaptive driving assistant model (ADAM) for advising drivers of autonomous vehicles. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(3): 21:1–21:28, September 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3545994.

Hepenstal:2021:DCA

[HZKW21] Sam Hepenstal, Leishi Zhang, Neesha Kodagoda, and B. L. William Wong. Developing conversational agents for use in criminal investigations. ACM Transactions on Interactive Intelligent Systems (TIIS), 11 (3–4):25:1–25:35, December 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3444369.

Ishii:2013:GAC

[INN13] Ryo Ishii, Yukiko I. Nakano, and Toyoaki Nishida. Gaze awareness in conversational agents: Estimating a user’s conversational engagement from eye gaze. ACM Transactions on Interactive Intelligent Systems (TIIS), 3(2): 11:1–11:??, July 2013. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Ishii:2016:PWW

[IOKY16a] Ryo Ishii, Kazuhiro Otsuka, Shiro Kumano, and Junji Yamato. Prediction of who will be the next speaker and when using gaze behavior in multiparty meetings. ACM Transactions on Interactive Intelligent Systems (TIIS), 6 (1):4:1–4:??, May 2016. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Ishii:2016:URP

[IOKY16b] Ryo Ishii, Kazuhiro Otsuka, Shiro Kumano, and Junji Yamato. Using respiration to predict who will speak next and when in multiparty meetings. ACM Transactions on Interactive Intelligent Systems (TIIS), 6 (2):20:1–20:??, August 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Intharah:2019:HDI

[ITB19] Thanapong Intharah, Daniyar Turmukhambetov, and Gabriel J. Brostow. HILC: Domain-independent PbD system via computer vision and follow-up questions. ACM Transactions on Interactive Intelligent Systems (TIIS), 9(2–3):16:1– 16:??, April 2019. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3234508.

Jameson:2013:MJR

[Jam13] Anthony Jameson. In memoriam: John Riedl. ACM Transactions on Interactive Intelligent Systems (TIIS), 3 (3):13:1–13:??, October 2013. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Juvina:2019:TUT

[JCL^{+}19] Ion Juvina, Michael G. Collins, Othalia Larue, William G. Kennedy, Ewart De Visser, and Celso De Melo. Toward a unified theory of learned trust in interpersonal and human-machine interactions. ACM Transactions on Interactive Intelligent Systems (TIIS), 9(4):

24:1–24:??, December 2019. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3230735.

Jokinen:2013:GTT

[JFNY13] Kristiina Jokinen, Hirohisa Furukawa, Masafumi Nishida, and Seiichi Yamamoto. Gaze and turntaking behavior in casual conversational interactions. ACM Transactions on Interactive Intelligent Systems (TIIS), 3(2):12:1–12:??, July 2013. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Jang:2022:RAH

[JIKB22] Yi Hyun Jang, Soo Han Im, Younah Kang, and Joon Sang Baek. Relational agents for the homeless with tuberculosis experience: Providing social support through humanagent relationships. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(2):15:1–15:22, June 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3488056.

Jugovac:2017:IRO

[JJ17] Michael Jugovac and Dietmar Jannach. Interacting with recommenders-overview and research directions. ACM Transactions on Interactive Intelligent Systems (TIIS), 7 (3):10:1–10:??, October 2017. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Jannach:2016:SDM

[JJL16] Dietmar Jannach, Michael Jugovac, and Lukas Lerche. Supporting the design of machine learning workflows with a recommendation system. ACM Transactions on Interactive Intelligent Systems (TIIS), 6(1):8:1–8:??, May 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Jorge:2024:HSA

[JJT24] Carolina Centeio Jorge, Catholijn M. Jonker, and Myrthe L. Tielman. How should an AI trust its human teammates? Exploring possible cues of artificial trust. ACM Transactions on Interactive Intelligent Systems (TIIS), 14(1): 5:1–5:??, March 2024. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3635475.

Jentner:2023:VAC

[JLH^{+}23] Wolfgang Jentner, Giuliana Lindholz, Hanna Hauptmann, Mennatallah El-Assady, Kwan-Liu Ma, and Daniel Keim. Visual analytics of co-occurrences to discover subspaces in structured data. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(2): 10:1–10:??, June 2023. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3579031.

Jameson:2011:ITI

[JR11] Anthony Jameson and John Riedl. Introduction to the Transactions on Interactive Intelligent Systems. ACM Transactions on Interactive Intelligent Systems (TIIS), 1 (1):1:1–1:??, October 2011. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Jameson:2012:ISI

[JR12] Anthony Jameson and John Riedl. Introduction to the special issue on highlights of the decade in interactive intelligent systems. ACM Transactions on Interactive Intelligent Systems (TIIS), 2(4):19:1–19:??, December 2012. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Jawaheer:2014:MUP

[JWK14] Gawesh Jawaheer, Peter Weller, and Patty Kostkova. Modeling user preferences in recommender systems: a classification framework for explicit and implicit user feedback. ACM Transactions on Interactive Intelligent Systems (TIIS), 4(2): 8:1–8:??, July 2014. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Kveton:2016:MIC

[KB16] Branislav Kveton and Shlomo Berkovsky. Minimal interaction content discovery in recommender systems. ACM Transactions on Interactive Intelligent Systems (TIIS), 6 (2):15:1–15:??, August 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Kaminskas:2017:DSN

[KB17] Marius Kaminskas and Derek Bridge. Diversity, serendipity, novelty, and coverage: a survey and empirical analysis of beyond-accuracy objectives in recommender systems. ACM Transactions on Interactive Intelligent Systems (TIIS), 7(1):2:1– 2:??, March 2017. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Krokos:2019:EDL

[KCC^{+}19] Eric Krokos, Hsueh-Chen Cheng, Jessica Chang, Bohdan Nebesh, Celeste Lyn Paul, Kirsten Whitley, and Amitabh Varshney. Enhancing deep learning with visual interactions. ACM Transactions on Interactive Intelligent Systems (TIIS), 9(1):

5:1–5:??, February 2019. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Kruse:2022:EMA

[KCM22] Jan Kruse, Andy M. Connor, and Stefan Marks. Evaluation of a multi-agent “Human-in-the-loop” game design system. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(3): 19:1–19:26, September 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3531009.

Koh:2019:DHG

[KCTH19] Jung In Koh, Josh Cherian, Paul Taele, and Tracy Hammond. Developing a hand gesture recognition system for mapping symbolic hand gestures to analogous emojis in computer-mediated communication. ACM Transactions on Interactive Intelligent Systems (TIIS), 9(1): 6:1–6:??, February 2019. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Kulahcioglu:2020:AAW

[KD20] Tugba Kulahcioglu and Gerard De Melo. Affect-aware word clouds. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(4): 34:1–34:25, December 2020.

CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3370928.

Khanna:2022:FAF

[KDA^{+}22] Roli Khanna, Jonathan Dodge, Andrew Anderson, Rupika Dikkala, Jed Irvine, Zeyad Shureih, Kin-Ho Lam, Caleb R. Matthews, Zhengxian Lin, Minsuk Kahng, Alan Fern, and Margaret Burnett. Finding AI’s faults with AAR/AI: an empirical study. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(1): 1:1–1:33, March 2022. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3487065.

Kaptein:2012:APS

[KDMA12] Maurits Kaptein, Boris De Ruyter, Panos Markopoulos, and Emile Aarts. Adaptive persuasive systems: a study of tailored persuasive text messages to reduce snacking. ACM Transactions on Interactive Intelligent Systems (TIIS), 2(2):10:1–10:??, June 2012. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Keizer:2014:MLS

[KFWL14] Simon Keizer, Mary Ellen Foster, Zhuoran Wang, and Oliver Lemon. Machine learning for social multiparty human–robot interaction. ACM Transactions on Interactive Intelligent Systems (TIIS), 4(3):14:1–14:??, October 2014. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Knott:2018:ATI

[KGCC18] Benjamin A. Knott, Jonathan Gratch, Angelo Cangelosi, and James Caverlee. ACM Transactions on Interactive Intelligent Systems (TiiS) special issue on trust and influence in intelligent humanmachine interaction. ACM Transactions on Interactive Intelligent Systems (TIIS), 8(4):25:1–25:??, November 2018. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3281451.

Kuo:2012:PRM

[KH12] Yen-Ling Kuo and Jane Yung-Jen Hsu. Planning for reasoning with multiple common sense knowledge bases. ACM Transactions on Interactive Intelligent Systems (TIIS), 2(3):17:1–17:??, September 2012. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Kay:2012:CPS

[KK12] Judy Kay and Bob Kummerfeld. Creating personalized systems that people can scrutinize and control: Drivers, principles and experience. ACM Transactions on Interactive Intelligent Systems (TIIS), 2(4): 24:1–24:??, December 2012. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Knijnenburg:2013:MDA

[KK13] Bart P. Knijnenburg and Alfred Kobsa. Making decisions about privacy: Information disclosure in context-aware recommender systems. ACM Transactions on Interactive Intelligent Systems (TIIS), 3 (3):20:1–20:??, October 2013. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Katsuragawa:2019:BLT

[KKL^{+}19] Keiko Katsuragawa, Ankit Kamal, Qi Feng Liu, Matei Negulescu, and Edward Lank. Bi-level thresholding: Analyzing the effect of repeated errors in gesture input. ACM Transactions on Interactive Intelligent Systems (TIIS), 9(2–3): 15:1–15:??, April 2019. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3181672.

Kronenberg:2022:IOW

[KKS22] Rotem Kronenberg, Tsvi Kuflik, and Ilan Shimshoni.

Improving office workers’ workspace using a selfadjusting computer screen. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(3):24:1– 24:32, September 2022. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3545993.

Kim:2021:LGR

[KLC^{+}21] Chris Kim, Xiao Lin, Christopher Collins, Graham W. Taylor, and Mohamed R. Amer. Learn, generate, rank, explain: a case study of visual explanation by generative machine learning. ACM Transactions on Interactive Intelligent Systems (TIIS), 11(3–4):23:1–23:34, December 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3465407.

Kocaballi:2022:SIC

[KLC^{+}22] A. Baki Kocaballi, Liliana Laranjo, Leigh Clark, Rafał Kocielnik, Robert J. Moore, Q. Vera Liao, and Timothy Bickmore. Special issue on conversational agents for healthcare and wellbeing. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(2):9:1–9:3, June 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3532860.

Koskela:2018:PIR

[KLR^{+}18] Markus Koskela, Petri Luukkonen, Tuukka Ruotsalo, Mats SjÖberg, and Patrik Floréen. Proactive information retrieval by capturing search intent from primary task context. ACM Transactions on Interactive Intelligent Systems (TIIS), 8(3):20:1–20:??, August 2018. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Khan:2020:PUM

[KMAM20] Euna Mehnaz Khan, Md. Saddam Hossain Mukta, Mohammed Eunus Ali, and Jalal Mahmud. Predicting users’ movie preference and rating behavior from personality and values. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(3):22:1–22:25, November 2020. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3338244.

Koulouri:2022:CSY

[KMO22] Theodora Koulouri, Robert D. Macredie, and David Olakitan. Chatbots to support young adults’ mental health: an exploratory study of acceptability. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(2):

11:1–11:39, June 2022. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3485874.

Kim:2018:HLS

[KP18] Bongjun Kim and Bryan Pardo. A human-in-the-loop system for sound event detection and annotation. ACM Transactions on Interactive Intelligent Systems (TIIS), 8(2):13:1–13:??, July 2018. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Karimi:2022:TTS

[KPMHB22] Pegah Karimi, Emanuele Plebani, Aqueasha Martin-Hammond, and Davide Bolchini. Textflow: Toward supporting screen-free manipulation of situation-relevant smart messages. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(4): 31:1–31:??, December 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3519263.

Kritikos:2014:TMD

[KPP14] K. Kritikos, D. Plexousakis, and F. Paternò. Task modeldriven realization of interactive application functionality through services. ACM Transactions on Interactive Intelligent Systems (TIIS), 3 (4):25:1–25:??, January 2014. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Kucher:2017:ALV

[KPSK17] Kostiantyn Kucher, Carita Paradis, Magnus Sahlgren, and Andreas Kerren. Active learning and visual analytics for stance classification with ALVA. ACM Transactions on Interactive Intelligent Systems (TIIS), 7 (3):14:1–14:??, October 2017. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Kumar:2014:TES

[KR14] Rohit Kumar and Carolyn P. Rosé. Triggering effective social support for online groups. ACM Transactions on Interactive Intelligent Systems (TIIS), 3 (4):24:1–24:??, January 2014. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Kouki:2020:GUP

[KSP^{+}20] Pigi Kouki, James Schaffer, Jay Pujara, John O’Donovan, and Lise Getoor. Generating and understanding personalized explanations in hybrid recommender systems. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(4):31:1–31:40, December 2020. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3365843.

Kulesza:2011:WOE

[KSW^{+}11] Todd Kulesza, Simone Stumpf, Weng-Keen Wong, Margaret M. Burnett, Stephen Perona, Andrew Ko, and Ian Oberst. Why-oriented end-user debugging of naive Bayes text classification. ACM Transactions on Interactive Intelligent Systems (TIIS), 1(1):2:1–2:??, October 2011. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Kim:2017:MML

[KTD17] Seungjun Kim, Dan Tasse, and Anind K. Dey. Making machine-learning applications for time-series sensor data graphical and interactive. ACM Transactions on Interactive Intelligent Systems (TIIS), 7(2): 8:1–8:??, July 2017. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Knijnenburg:2016:ICI

[KW16] Bart P. Knijnenburg and Martijn C. Willemsen. Inferring capabilities of intelligent agents from their external traits. ACM Transactions on Interactive Intelligent Systems (TIIS), 6(4): 28:1–28:??, December 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Li:2021:ASE

[LBC^{+}21] Mingzhao Li, Zhifeng Bao, Farhana Choudhury, Hanan Samet, Matt Duckham, and Timos Sellis. AOI-shapes: an efficient footprint algorithm to support visualization of user-defined urban areas of interest. ACM Transactions on Interactive Intelligent Systems (TIIS), 11 (3–4):27:1–27:32, December 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3431817.

Lin:2018:GES

[LC18] Yu-Ru Lin and Nan Cao. Guest editorial: Special issue on interactive visual analysis of human and crowd behaviors. ACM Transactions on Interactive Intelligent Systems (TIIS), 8(1): 1:1–1:??, March 2018. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Liu:2023:SGL

[LCKM23] Huimin Liu, Minsoo Choi, Dominic Kao, and Christos Mousas. Synthesizing game levels for collaborative gameplay in a shared virtual environment. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(1): 2:1–2:??, March 2023. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3558773.

Lee:2023:LAE

[LDLW23] Benjamin Charles Germain Lee, Doug Downey, Kyle Lo, and Daniel S. Weld. LIMEADE: From AI explanations to advice taking. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(4):24:1–24:??, December 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL

https://dl.acm.org/doi/
10.1145/3589345.

Larasati:2023:MEE

[LDM23] Retno Larasati, Anna De Liddo, and Enrico Motta. Meaningful explanation effect on User’s trust in an AI medical system: Designing explanations for nonexpert users. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(4): 30:1–30:??, December 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3631614.

Liu:2022:SSE

[LDS^{+}22] Fang Liu, Xiaoming Deng, Jiancheng Song, Yu-Kun Lai, Yong-Jin Liu, Hao Wang, Cuixia Ma, Shengfeng Qin, and Hongan Wang. SketchMaker: Sketch extraction and reuse for interactive scene sketch composition. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(3): 23:1–23:26, September 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3543956.

Lieberman:2012:ISI

[LH12] Henry Lieberman and Catherine Havasi. Introduction to the Special Issue on Common Sense for Interactive Systems. ACM Transactions on Interactive Intelligent Systems (TIIS), 2(3):14:1–14:??, September 2012. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Lin:2023:PIM

[LL23] Yi-Ling Lin and Shao-Wei Lee. A personalized interaction mechanism framework for micro-moment recommender systems. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(1):4:1–4:??, March 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3569586.

Lee:2019:UAM

[LPBG19] Junghyo Lee, Prajwal Paudyal, Ayan Banerjee, and Sandeep K. S. Gupta. A user-adaptive modeling for eating action identification from wristband time series. ACM Transactions on Interactive Intelligent Systems (TIIS), 9(4): 22:1–22:??, December 2019. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3300149.

Li:2023:VAN

[LWFM23] Yiran Li, Junpeng Wang, Takanori Fujiwara, and Kwan-Liu Ma. Visual analytics of neuron vulnerability to adversarial attacks on convolutional neural networks. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(4):20:1–20:??, December 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL

https://dl.acm.org/doi/
10.1145/3587470.

Li:2023:ESE

[LZL^{+}23] Xingjun Li, Yizhi Zhang, Justin Leung, Chengnian Sun, and Jian Zhao. EDAssistant: Supporting exploratory data analysis in computational notebooks with in situ code search and recommendation. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(1):1:1–1:??, March 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3545995.

Liu:2024:SCM

[LZW^{+}24] Can Liu, Yu Zhang, Cong Wu, Chen Li, and Xiaoru Yuan. A spatial constraint model for manipulating static visualizations. ACM Transactions on Interactive Intelligent Systems (TIIS), 14(2): 12:1–12:??, June 2024. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3657642.

Martens:2014:ISI

[Mar14] Jean-Bernard Martens. Interactive statistics with Illmo. ACM Transactions on Interactive Intelligent Systems (TIIS), 4(1):4:1–4:??, April 2014. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Mazilu:2015:WAG

[MBD^{+}15] Sinziana Mazilu, Ulf Blanke, Moran Dorfman, Eran Gazit, Anat Mirelman, Jeffrey M. Hausdorff, and Gerhard Tröster. A wearable assistant for gait training for Parkinson’s disease with freezing of gait in out-of-the-lab environments. ACM Transactions on Interactive Intelligent Systems (TIIS), 5(1): 5:1–5:??, March 2015. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Mahmoud:2016:AAN

[MBR16] Marwa Mahmoud, Tadas Baltrusaitis, and Peter Robinson. Automatic analysis of naturalistic hand-over-face gestures. ACM Transactions on Interactive Intelligent Systems (TIIS), 6 (2):19:1–19:??, August 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Mo:2024:CMO

[MDC^{+}24] George Mo, John Dudley, Liwei Chan, Yi-Chi Liao, Antti Oulasvirta, and Per Ola Kristensson. Cooperative multiobjective Bayesian design optimization. ACM Transactions on Interactive Intelligent Systems (TIIS), 14(2): 13:1–13:??, June 2024. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3657643.

Maharjan:2022:ESE

[MDR^{+}22] Raju Maharjan, Kevin Doherty, Darius Adam Rohani, Per Bækgaard, and Jakob E. Bardram. Experiences of a speech-enabled conversational agent for the selfreport of well-being among people living with affective disorders: an in-the-wild study. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(2): 10:1–10:29, June 2022. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3484508.

Molyneaux:2013:CAM

[MGF13] David Molyneaux, Hans Gellersen, and Joe Finney. Cooperative augmentation of mobile smart objects with projected displays. ACM Transactions on Interactive Intelligent Systems (TIIS), 3 (2):7:1–7:??, July 2013. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Mascarenhas:2022:FTT

[MGP^{+}22] Samuel Mascarenhas, Manuel Guimarães, Rui Prada, Pedro A. Santos, João Dias, and Ana Paiva. FAtiMA toolkit: Toward an accessible tool for the development of socio-emotional agents. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(1):8:1–8:30, March 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3510822.

Morrison:2018:VUS

[MHC^{+}18] Cecily Morrison, Kit Huckvale, Bob Corish, Richard Banks, Martin Grayson, Jonas Dorn, Abigail Sellen, and Sân Lindley. Visualizing ubiquitously sensed measures of motor ability in multiple sclerosis: Reflections on communicating machine learning in practice. ACM Transactions on Interactive Intelligent Systems (TIIS), 8(2): 12:1–12:??, July 2018. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Mehrotra:2024:IBE

[MJJT24] Siddharth Mehrotra, Carolina Centeio Jorge, Catholijn M. Jonker, and Myrthe L. Tielman. Integrity-based explanations for fostering appropriate trust in AI agents. ACM Transactions on Interactive Intelligent Systems (TIIS), 14(1):4:1–4:??, March 2024. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL

https://dl.acm.org/doi/
10.1145/3610578.

Meignan:2015:RTI

[MKF^{+}15] David Meignan, Sigrid Knust, Jean-Marc Frayret, Gilles Pesant, and Nicolas Gaud. A review and taxonomy of interactive optimization methods in operations research. ACM Transactions on Interactive Intelligent Systems (TIIS), 5 (3):17:1–17:??, October 2015. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Masai:2017:EFE

[MKS^{+}17] Katsutoshi Masai, Kai Kunze, Yuta Sugiura, Masa Ogata, Masahiko Inami, and Maki Sugimoto. Evaluation of facial expression recognition by a smart eyewear for facial direction changes, repeatability, and positional drift. ACM Transactions on Interactive Intelligent Systems (TIIS), 7(4):15:1–15:??, December 2017. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Mousas:2021:PSV

[MKW21] Christos Mousas, Claudia Krogmeier, and Zhiquan Wang. Photo sequences of varying emotion: Optimization with a valence-arousal annotated dataset. ACM Transactions on Interactive Intelligent Systems (TIIS), 11(2):16:1–16:19, July 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3458844.

Mihoub:2019:WSS

[ML19] Alaeddine Mihoub and Grégoire Lefebvre. Wearables and social signal processing for smarter public presentations. ACM Transactions on Interactive Intelligent Systems (TIIS), 9(2–3):9:1–9:??, April 2019. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3234507.

Ma:2021:HTR

[MLD^{+}21] Wanqi Ma, Xiaoxiao Liao, Wei Dai, Weike Pan, and Zhong Ming. Holistic transfer to rank for top-N recommendation. ACM Transactions on Interactive Intelligent Systems (TIIS), 11(1): 8:1, April 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL

https://dl.acm.org/doi/
10.1145/3434360.

Martinez:2023:DEH

[MNL^{+}23] Miguel Angel Meza Martínez, Mario Nadj, Moritz Langner, Peyman Toreini, and Alexander Maedche. Does this explanation help? Designing local model-agnostic explanation representations and an experimental evaluation using eye-tracking technology. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(4):27:1–27:??, December 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL

https://dl.acm.org/doi/
10.1145/3607145.

Mohan:2021:ERC

[Moh21] Shiwali Mohan. Exploring the role of common model of cognition in designing adaptive coaching interactions for health behavior change. ACM Transactions on Interactive Intelligent Systems (TIIS), 11(1):1:1–1:30, April 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3375790.

Marge:2019:MDR

[MR19] Matthew Marge and Alexander I. Rudnicky. Miscommunication detection and recovery in situated human– robot dialogue. ACM Transactions on Interactive Intelligent Systems (TIIS), 9(1): 3:1–3:??, February 2019. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3237189.

Malik:2016:HVH

[MSD^{+}16] Sana Malik, Ben Shneiderman, Fan Du, Catherine Plaisant, and Margret Bjarnadottir. High-volume hypothesis testing: Systematic exploration of event sequence comparisons. ACM Transactions on Interactive Intelligent Systems (TIIS), 6(1): 9:1–9:??, May 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Mohanty:2020:PSI

[MTML20] Vikram Mohanty, David Thames, Sneha Mehta, and Kurt Luther. Photo Sleuth: Identifying historical portraits with face recognition and crowdsourced human expertise. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(4):

33:1–33:36, December 2020. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3365842.

Mohan:2020:DAH

[MVH20] Shiwali Mohan, Anusha Venkatakrishnan, and Andrea L. Hartzler. Designing an AI health coach and studying its utility in promoting regular aerobic exercise. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(2): 14:1–14:30, June 2020. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3366501.

Mutlu:2016:VRP

[MVT16] Belgin Mutlu, Eduardo Veas, and Christoph Trattner. VizRec: Recommending personalized visualizations. ACM Transactions on Interactive Intelligent Systems (TIIS), 6(4):31:1–31:??, December 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Meng:2021:VVA

[MWP^{+}21] Linhao Meng, Yating Wei, Rusheng Pan, Shuyue Zhou, Jianwei Zhang, and Wei Chen. VADAF: Visualization for abnormal client detection and analysis in federated learning. ACM Transactions on Interactive Intelligent Systems (TIIS), 11 (3–4):26:1–26:23, December 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3426866.

Mohseni:2021:MSF

[MZR21] Sina Mohseni, Niloofar Zarei, and Eric D. Ragan. A multidisciplinary survey and framework for design and evaluation of explainable AI systems. ACM Transactions on Interactive Intelligent Systems (TIIS), 11 (3–4):24:1–24:45, December 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3387166.

Nakano:2016:ISI

[NBHJ16] Yukiko I. Nakano, Roman Bednarik, Hung-Hsuan Huang, and Kristiina Jokinen. Introduction to the special issue on new directions in eye gaze for interactive intelligent systems. ACM Transactions on Interactive Intelligent Systems (TIIS), 6(1): 1:1–1:??, May 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Nourbakhsh:2017:DUC

[NCWC17] Nargess Nourbakhsh, Fang Chen, Yang Wang, and Rafael A. Calvo. Detecting users’ cognitive load by galvanic skin response with affective interference. ACM Transactions on Interactive Intelligent Systems (TIIS), 7 (3):12:1–12:??, October 2017. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Niewiadomski:2019:AMQ

[NKP^{+}19] Radoslaw Niewiadomski, Ksenia Kolykhalova, Stefano Piana, Paolo Alborno, Gualtiero Volpe, and Antonio Camurri. Analysis of movement quality in full-body physical activities. ACM Transactions on Interactive Intelligent Systems (TIIS), 9(1): 1:1–1:??, February 2019. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3132369.

Ngo:2014:EIM

[NLN^{+}14] Hung Ngo, Matthew Luciw, Jawas Nagi, Alexander Forster, Jürgen Schmidhuber, and Ngo Anh Vien. Efficient interactive multiclass learning from binary feedback. ACM Transactions on Interactive Intelligent Systems (TIIS), 4(3): 12:1–12:??, August 2014. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Nourani:2022:IUB

[NRB^{+}22] Mahsan Nourani, Chiradeep Roy, Jeremy E. Block, Donald R. Honeycutt, Tahrima Rahman, Eric D. Ragan, and Vibhav Gogate. On the importance of user backgrounds and impressions: Lessons learned from interactive AI applications. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(4): 28:1–28:??, December 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3531066.

Nakao:2022:TIE

[NSA^{+}22] Yuri Nakao, Simone Stumpf, Subeida Ahmed, Aisha Naseer, and Lorenzo Strappelli. Toward involving end-users in interactive human-in-the-loop AI fairness. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(3):18:1–18:30, September 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3514258.

Nourashrafeddin:2018:VAI

[NSMM18] Seyednaser Nourashrafeddin, Ehsan Sherkat, Rosane Minghim, and Evangelos E. Milios. A visual approach for interactive keyterm-based clustering. ACM Transactions on Interactive Intelligent Systems (TIIS), 8(1): 6:1–6:??, March 2018. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Narzt:2018:ECA

[NWP^{+}18] Wolfgang Narzt, Otto Weichselbaum, Gustav Pomberger, Markus Hofmarcher, Michael Strauss, Peter Holzkorn, Roland Haring, and Monika Sturm. Estimating collective attention toward a public display. ACM Transactions on Interactive Intelligent Systems (TIIS), 8 (3):21:1–21:??, August 2018. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Nakano:2016:GRG

[NYYT16] Yukiko I. Nakano, Takashi Yoshino, Misato Yatsushiro, and Yutaka Takase. Generating robot gaze on the basis of participation roles and dominance estimation in multiparty interaction. ACM Transactions on Interactive Intelligent Systems (TIIS), 5 (4):22:1–22:??, January 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Oraby:2019:MCC

[OBG^{+}19] Shereen Oraby, Mansurul Bhuiyan, Pritam Gundecha, Jalal Mahmud, and Rama Akkiraju. Modeling and computational characterization of Twitter customer service conversations. ACM Transactions on Interactive Intelligent Systems (TIIS), 9(2–3): 18:1–18:??, April 2019. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3213014.

Oviatt:2018:DHS

[OHZ^{+}18] S. Oviatt, K. Hang, J. Zhou, K. Yu, and F. Chen. Dynamic handwriting signal features predict domain expertise. ACM Transactions on Interactive Intelligent Systems (TIIS), 8(3):18:1–18:??, August 2018. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Oviatt:2021:KWY

[OLS21] Sharon Oviatt, Jionghao Lin, and Abishek Sriramulu. I know what you know: What hand movements reveal about domain expertise. ACM Transactions on Interactive Intelligent Systems (TIIS), 11(1):4:1–4:26, April 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3423049.

Okita:2011:MAA

[ONTHS11] Sandra Y. Okita, Victor Ng-Thow-Hing, and Ravi K. Sarvadevabhatla. Multimodal approach to affective humanrobot interaction design with children. ACM Transactions on Interactive Intelligent Systems (TIIS), 1(1): 5:1–5:??, October 2011. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Ochs:2017:UPB

[OPM17] Magalie Ochs, Catherine Pelachaud, and Gary Mckeown. A user perception– based approach to create smiling embodied conversational agents. ACM Transactions on Interactive Intelligent Systems (TIIS), 7(1): 4:1–4:??, March 2017. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Pejsa:2015:GAM

[PAGM15] Tomislav Pejsa, Sean Andrist, Michael Gleicher, and Bilge Mutlu. Gaze and attention management for embodied conversational agents. ACM Transactions on Interactive Intelligent Systems (TIIS), 5(1):3:1–3:??, March 2015. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Prendinger:2016:IBT

[PASR^{+}16] Helmut Prendinger, Nahum Alvarez, Antonio Sanchez-Ruiz, Marc Cavazza, joão Catarino, João Oliveira, Rui Prada, Shuji Fujimoto, and Mika Shigematsu. Intelligent biohazard training based on real-time task recognition. ACM Transactions on Interactive Intelligent Systems (TIIS), 6(3):21:1–21:??, October 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Panigutti:2023:CDH

[PBF^{+}23] Cecilia Panigutti, Andrea Beretta, Daniele Fadda, Fosca Giannotti, Dino Pedreschi, Alan Perotti, and Salvatore Rinzivillo. Codesign of human-centered, explainable AI for clinical decision support. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(4): 21:1–21:??, December 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3587271.

Pan:2020:SID

[PBKS20] Shimei Pan, Oliver Brdiczka, Andrea Kleinsmith, and Yangqiu Song. Special issue on data-driven personality modeling for intelligent human-computer interaction. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(3):17:1–17:3, November 2020. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL

https://dl.acm.org/doi/
10.1145/3402522.

Poesio:2013:PDU

[PCK^{+}13] Massimo Poesio, Jon Chamberlain, Udo Kruschwitz, Livio Robaldo, and Luca Ducceschi. Phrase detectives: Utilizing collective intelligence for Internet-scale language resource creation. ACM Transactions on Interactive Intelligent Systems (TIIS), 3(1):3:1–3:??, April 2013. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Polack:2018:CIM

[PCK^{+}18] Peter J. Polack Jr., Shang-Tse Chen, Minsuk Kahng, Kaya De Barbaro, Rahul Basole, Moushumi Sharmin, and Duen Horng Chau. Chronodes: Interactive multifocus exploration of event sequences. ACM Transactions on Interactive Intelligent Systems (TIIS), 8(1): 2:1–2:??, March 2018. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Paudel:2017:UAD

[PCNB17] Bibek Paudel, Fabian Christoffel, Chris Newell, and Abraham Bernstein. Updatable, accurate, diverse, and scalable recommendations for interactive applications. ACM Transactions on Interactive Intelligent Systems (TIIS), 7(1):1:1–1:??, March 2017. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Penney:2021:SGE

[PDA^{+}21] Sean Penney, Jonathan Dodge, Andrew Anderson, Claudia Hilderbrand, Logan Simpson, and Margaret Burnett. The shoutcasters, the game enthusiasts, and the AI: Foraging for explanations of real-time strategy players. ACM Transactions on Interactive Intelligent Systems (TIIS), 11(1):2:1–2:46, April 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3396047.

Park:2012:CFM

[PKCS12] Souneil Park, Seungwoo Kang, Sangyoung Chung, and Junehwa Song. A computational framework for media bias mitigation. ACM Transactions on Interactive Intelligent Systems (TIIS), 2(2): 8:1–8:??, June 2012. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Paudyal:2019:CTS

[PLBG19] Prajwal Paudyal, Junghyo Lee, Ayan Banerjee, and Sandeep K. S. Gupta. A comparison of techniques for sign language alphabet recognition using armband wearables. ACM Transactions on Interactive Intelligent Systems (TIIS), 9(2–3):14:1– 14:??, April 2019. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3150974.

Paiva:2017:EVA

[PLBW17] Ana Paiva, Iolanda Leite, Hana Boukricha, and Ipke Wachsmuth. Empathy in virtual agents and robots: a survey. ACM Transactions on Interactive Intelligent Systems (TIIS), 7 (3):11:1–11:??, October 2017. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Park:2016:MAP

[PSC^{+}16] Sunghyun Park, Han Suk Shim, Moitreya Chatterjee, Kenji Sagae, and Louis-Philippe Morency. Multimodal analysis and prediction of persuasiveness in online social multimedia. ACM Transactions on Interactive Intelligent Systems (TIIS), 6 (3):25:1–25:??, October 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Park:2018:MFU

[PSMO18] Souneil Park, Joan Serrà, Enrique Frias Martinez, and Nuria Oliver. MobInsight: a framework using semantic neighborhood features for localized interpretations of urban mobility. ACM Transactions on Interactive Intelligent Systems (TIIS), 8 (3):23:1–23:??, August 2018.

CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Piana:2016:ABG

[PSOC16] Stefano Piana, Alessandra Staglianò, Francesca Odone, and Antonio Camurri. Adaptive body gesture representation for automatic emotion recognition. ACM Transactions on Interactive Intelligent Systems (TIIS), 6(1): 6:1–6:??, May 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Pham:2019:AMA

[PW19] Phuong Pham and Jingtao Wang. AttentiveVideo: a multimodal approach to quantify emotional responses to mobile advertisements. ACM Transactions on Interactive Intelligent Systems (TIIS), 9(2–3):8:1–8:??, April 2019. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3232233.

Pan:2016:TLS

[PYDM16] Weike Pan, Qiang Yang, Yuchao Duan, and Zhong Ming. Transfer learning for semisupervised collaborative recommendation. ACM Transactions on Interactive Intelligent Systems (TIIS), 6 (2):10:1–10:??, August 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Rafailidis:2014:CBT

[RAE^{+}14] Dimitrios Rafailidis, Apostolos Axenopoulos, Jonas Etzold, Stavroula Manolopoulou, and Petros Daras. Contentbased tag propagation and tensor factorization for personalized item recommendation based on social tagging. ACM Transactions on Interactive Intelligent Systems (TIIS), 3(4):26:1–26:??, January 2014. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Ramirez-Amaro:2017:AVG

[RAMZ^{+}17]  Karinne  Ramirez-Amaro,

Humera Noor Minhas, Michael Zehetleitner, Michael Beetz, and Gordon Cheng. Added value of gaze-exploiting semantic representation to allow robots inferring human behaviors. ACM Transactions on Interactive Intelligent Systems (TIIS), 7(1): 5:1–5:??, March 2017. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Roffarello:2021:UDM

[RD21] Alberto Monge Roffarello and Luigi De Russis. Understanding, discovering, and mitigating habitual smartphone use in young adults. ACM Transactions on Interactive Intelligent Systems (TIIS), 11(2):13:1– 13:34, July 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL

https://dl.acm.org/doi/
10.1145/3447991.

Rathore:2024:VVI

[RDP^{+}24] Archit Rathore, Sunipa Dev, Jeff M. Phillips, Vivek Srikumar, Yan Zheng, Chin-Chia Michael Yeh, Junpeng Wang, Wei Zhang, and Bei Wang. VERB: Visualizing and interpreting bias mitigation techniques geometrically for word representations. ACM Transactions on Interactive Intelligent Systems (TIIS), 14(1): 3:1–3:??, March 2024. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3604433.

Radeta:2024:MME

[RFR^{+}24] Marko Radeta, Ruben Freitas, Claudio Rodrigues, Agustin Zuniga, Ngoc Thi Nguyen, Huber Flores, and Petteri Nurmi. Man and the machine: Effects of AI-assisted human labeling on interactive annotation of real-time video streams. ACM Transactions on Interactive Intelligent Systems (TIIS), 14(2): 10:1–10:??, June 2024. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3649457.

Ramachandran:2019:TER

[RHS19] Aditi Ramachandran, Chien-Ming Huang, and Brian Scassellati. Toward effective robot–child tutoring: Internal motivation, behavioral intervention, and learning outcomes. ACM Transactions on Interactive Intelligent Systems (TIIS), 9(1): 2:1–2:??, February 2019. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3213768.

Riveiro:2014:ENM

[Riv14] Maria Riveiro. Evaluation of normal model visualization for anomaly detection in maritime traffic. ACM Transactions on Interactive Intelligent Systems (TIIS), 4(1): 5:1–5:??, April 2014. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Rosenfeld:2016:PAD

[RK16] Ariel Rosenfeld and Sarit Kraus. Providing arguments in discussions on the basis of the prediction of human argumentative behavior. ACM Transactions on Interactive Intelligent Systems (TIIS), 6(4):30:1–30:??, December 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Roy:2023:EAR

[RNA^{+}23] Chiradeep Roy, Mahsan Nourani, Shivvrat Arya, Mahesh Shanbhag, Tahrima Rahman, Eric D. Ragan, Nicholas Ruozzi, and Vibhav Gogate. Explainable activity recognition in videos using deep learning and tractable probabilistic models. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(4):29:1–29:??, December 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3626961.

Rosenberg:2021:ECA

[RPRK^{+}21] Maor Rosenberg, Hae Won Park, Rinat Rosenberg-Kima, Safinah Ali, Anastasia K. Ostrowski, Cynthia Breazeal, and Goren Gordon. Expressive cognitive architecture for a curious social robot. ACM Transactions on Interactive Intelligent Systems (TIIS), 11(2): 12:1–12:25, July 2021. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3451531.

Ramos:2022:FAO

[RRS^{+}22] Gonzalo Ramos, Napol Rachatasumrit, Jina Suh, Rachel Ng, and Christopher Meek. ForSense: Accelerating online research through sensemaking integration and machine research support. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(4):30:1–30:??, December 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL

https://dl.acm.org/doi/
10.1145/3532853.

Rudrauf:2023:CPC

[RSPT^{+}23] D. Rudrauf, G. Sergeant-Perhtuis, Y. Tisserand, T. Monnor, V. De Gevigney, and O. Belli. Combining the projective consciousness model and virtual humans for immersive psychological research: a proof-of-concept simulating a ToM assessment. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(2): 8:1–8:??, June 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL

https://dl.acm.org/doi/
10.1145/3583886.

Razavi:2022:DBO

[RSvO^{+}22] S. Zahra Razavi, Lenhart K.

Schubert, Kimberly van Orden, Mohammad Rafayet Ali, Benjamin Kane, and Ehsan Hoque. Discourse behavior of older adults interacting with a dialogue agent competent in multiple topics. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(2):14:1–14:21, June 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3484510.

Segura:2021:BBO

[SB21] Vinícius Segura and Simone D. J. Barbosa. BONNIE: Building online narratives from noteworthy interaction events. ACM Transactions on Interactive Intelligent Systems (TIIS), 11 (3–4):21:1–21:31, December 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3423048.

Sidner:2018:CNT

[SBN^{+}18] Candace L. Sidner, Timothy Bickmore, Bahador Nooraie, Charles Rich, Lazlo Ring, Mahni Shayganfar, and Laura Vardoulakis. Creating new technologies for companionable agents to support isolated older adults. ACM Transactions on Interactive Intelligent Systems (TIIS), 8 (3):17:1–17:??, August 2018. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Steichen:2014:IVT

[SCC14] Ben Steichen, Cristina Conati, and Giuseppe Carenini. Inferring visualization task properties, user performance, and user cognitive abilities from eye gaze data. ACM Transactions on Interactive Intelligent Systems (TIIS), 4(2):11:1–11:??, July 2014. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Sanchez-Cortes:2015:MVM

[SCKOGP15] Dairazalia Sanchez-Cortes, Shiro Kumano, Kazuhiro Otsuka, and Daniel Gatica-Perez. In the mood for vlog: Multimodal inference in conversational social video. ACM Transactions on Interactive Intelligent Systems (TIIS), 5(2):9:1–9:??, July 2015. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Shatilov:2021:EEB

[SCLH21] Kirill A. Shatilov, Dimitris Chatzopoulos, Lik-Hang Lee, and Pan Hui. Emerging ExG-based NUI inputs in extended realities: a bottomup survey. ACM Transactions on Interactive Intelligent Systems (TIIS), 11(2): 10:1–10:49, July 2021. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3457950.

Smith:2023:GDB

[SD23] Ronnie Smith and Mauro Dragone. Generalisable dialogue-based approach for active learning of activities of daily living. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(3):

18:1–18:??, September 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3616017.

Song:2012:CBH

[SDD12] Yale Song, David Demirdjian, and Randall Davis. Continuous body and hand gesture recognition for natural human-computer interaction. ACM Transactions on Interactive Intelligent Systems (TIIS), 2(1):5:1–5:??, March 2012. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Self:2018:OLP

[SDW^{+}18] Jessica Zeitz Self, Michelle Dowling, John Wenskovitch, Ian Crandell, Ming Wang, Leanna House, Scotland Leman, and Chris North. Observation-level and parametric interaction for highdimensional data analysis. ACM Transactions on Interactive Intelligent Systems (TIIS), 8(2):15:1–15:??, July 2018. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Schrills:2023:HDU

[SF23] Tim Schrills and Thomas Franke. How do users experience traceability of AI systems? Examining subjective information processing awareness in automated insulin delivery (AID) systems.

ACM Transactions on Interactive Intelligent Systems (TIIS), 13(4):25:1–25:??, December 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL

https://dl.acm.org/doi/
10.1145/3588594.

Song:2020:FLT

[SFKL20] Jean Y. Song, Raymond Fok, Juho Kim, and Walter S. Lasecki. FourEyes: Leveraging tool diversity as a means to improve aggregate accuracy in crowdsourcing. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(1):3:1–3:30, January 2020. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3237188.

Song:2013:PII

[SFTI^{+}13] Wei Song, Andrew Finch, Kumiko Tanaka-Ishii, Keiji Yasuda, and Eiichiro Sumita. picoTrans: an intelligent icon-driven interface for cross-lingual communication. ACM Transactions on Interactive Intelligent Systems (TIIS), 3(1):5:1–5:??, April 2013. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Swanson:2012:SAU

[SG12] Reid Swanson and Andrew S.

Gordon. Say anything: Using textual case-based reasoning to enable open-domain interactive storytelling. ACM Transactions on Interactive Intelligent Systems (TIIS), 2(3):16:1–16:??, September 2012. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Suh:2020:AFS

[SGR^{+}20] Jina Suh, Soroush Ghorashi, Gonzalo Ramos, Nan-Chen Chen, Steven Drucker, Johan Verwey, and Patrice Simard. AnchorViz: Facilitating semantic data exploration and concept discovery for interactive machine learning. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(1):7:1–7:38, January 2020. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3241379.

Scherer:2012:SLN

[SGS^{+}12] Stefan Scherer, Michael Glodek, Friedhelm Schwenker, Nick Campbell, and Günther Palm. Spotting laughter in natural multiparty conversations: a comparison of automatic online and offline approaches using audiovisual data. ACM Transactions on Interactive Intelligent Systems (TIIS), 2(1): 4:1–4:??, March 2012. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Salah:2015:BUA

[SHA^{+}15a] Albert Ali Salah, Hayley Hung, Oya Aran, Hatice Gunes, and Matthew Turk. Behavior understanding for arts and entertainment. ACM Transactions on Interactive Intelligent Systems (TIIS), 5 (3):12:1–12:??, October 2015. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Salah:2015:BIS

[SHA^{+}15b] Albert Ali Salah, Hayley Hung, Oya Aran, Hatice Gunes, and Matthew Turk. Brief introduction to the special issue on behavior understanding for arts and entertainment. ACM Transactions on Interactive Intelligent Systems (TIIS), 5(2): 6:1–6:??, July 2015. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Sharma:2019:LSI

[SHK19] Mohit Sharma, F. Maxwell Harper, and George Karypis. Learning from sets of items in recommender systems. ACM Transactions on Interactive Intelligent Systems (TIIS), 9(4):19:1–19:??, December 2019. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3326128.

Santos:2019:PPT

[SHKM19] Carlos Pereira Santos, Kevin Hutchinson, Vassilis-Javed Khan, and Panos Markopoulos. Profiling personality traits with games. ACM Transactions on Interactive Intelligent Systems (TIIS), 9 (2–3):11:1–11:??, April 2019. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3230738.

Shneiderman:2020:BGB

[Shn20] Ben Shneiderman. Bridging the gap between ethics and practice: Guidelines for reliable, safe, and trustworthy human-centered AI systems. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(4): 26:1–26:31, December 2020. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3419764.

Sutherland:2016:EAE

[SHY16] Steven C. Sutherland, Casper Harteveld, and Michael E. Young. Effects of the advisor and environment on requesting and complying with automated advice. ACM Transactions on Interactive Intelligent Systems (TIIS), 6(4): 27:1–27:??, December 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Sevastjanova:2021:QGA

[SJS^{+}21] Rita Sevastjanova, Wolfgang Jentner, Fabian Sperrle, Rebecca Kehlbeck, Jürgen Bernard, and Mennatallah El-assady. QuestionComb: a gamification approach for the visual explanation of linguistic phenomena through interactive labeling. ACM Transactions on Interactive Intelligent Systems (TIIS), 11(3–4):19:1–19:38, December 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3429448.

Steptoe:2018:VAF

[SKG^{+}18] Michael Steptoe, Robert Krüger, Rolando Garcia, Xing Liang, and Ross Maciejewski. A visual analytics framework for exploring theme park dynamics. ACM Transactions on Interactive Intelligent Systems (TIIS), 8(1):4:1–4:??, March 2018. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Soto:2015:EVA

[SKKM15] Axel J. Soto, Ryan Kiros, Vlado Keselj, and Evangelos Milios. Exploratory visual analysis and interactive pattern extraction from semi-structured data. ACM Transactions on Interactive Intelligent Systems (TIIS), 5 (3):16:1–16:??, October 2015. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Spinner:2024:TEG

[SKS^{+}24] Thilo Spinner, Rebecca Kehlbeck, Rita Sevastjanova, Tobias Stähle, Daniel A. Keim, Oliver Deussen, and Mennatallah El-Assady. [treeemoji]-generAItor: Tree-in-the-loop text generation for language model explainability and adaptation. ACM Transactions on Interactive Intelligent Systems (TIIS), 14(2):14:1–14:??, June 2024. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3652028.

Spiller:2021:PVS

[SLH^{+}21] Moritz Spiller, Ying-Hsang Liu, Md Zakir Hossain, Tom Gedeon, Julia Geissler, and Andreas Nürnberger. Predicting visual search task success from eye gaze data as a basis for user-adaptive information visualization systems. ACM Transactions on Interactive Intelligent Systems (TIIS), 11(2):14:1– 14:25, July 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL

https://dl.acm.org/doi/
10.1145/3446638.

Schreiber:2013:ISI

[SLM^{+}13] Daniel Schreiber, Kris Luyten, Max Mühlhäuser, Oliver Brdiczka, and Melanie Hartman. Introduction to the special issue on interaction with smart objects. ACM Transactions on Interactive Intelligent Systems (TIIS), 3(2): 6:1–6:??, July 2013. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Smith:2018:PUC

[SLMK18] Jim Smith, Phil Legg, Milos Matovic, and Kristofer Kinsey. Predicting user confidence during visual decision making. ACM Transactions on Interactive Intelligent Systems (TIIS), 8(2): 10:1–10:??, July 2018. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Sluyters:2023:RAR

[SLVV23] Arthur Sluÿters, Sébastien Lambot, Jean Vanderdonckt, and Radu-Daniel Vatavu. RadarSense: Accurate recognition of mid-air hand gestures with radar sensing and few training examples. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(3):16:1–16:??, September 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3589645.

Shibata:2023:CCS

[SMF^{+}23] Ryoichi Shibata, Shoya Matsumori, Yosuke Fukuchi, Tomoyuki Maekawa, Mitsuhiko Kimoto, and Michita Imai. Conversational context-sensitive ad generation with a few corequeries. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(3): 15:1–15:??, September 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3588578.

Singh:2023:DEA

[SML^{+}23] Ronal Singh, Tim Miller, Henrietta Lyons, Liz Sonenberg, Eduardo Velloso, Frank Vetere, Piers Howe, and Paul Dourish. Directive explanations for actionable explainability in machine learning applications. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(4): 23:1–23:??, December 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3579363.

Sherkat:2020:VAA

[SMM20] Ehsan Sherkat, Evangelos E. Milios, and Rosane Minghim. A visual analytics approach for interactive document clustering. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(1): 6:1–6:33, January 2020. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3241380.

Svikhnushina:2022:PMK

[SP22] Ekaterina Svikhnushina and Pearl Pu. PEACE: a model of key social and emotional qualities of conversational chatbots. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(4): 32:1–32:??, December 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3531064.

Sen:2019:TUS

[SSL^{+}19] Shilad Sen, Anja Beth Swoap, Qisheng Li, Ilse Dippenaar, Monica Ngo, Sarah Pujol, Rebecca Gold, Brooke Boatman, Brent Hecht, and Bret Jackson. Toward universal spatialization through Wikipedia-based semantic enhancement. ACM Transactions on Interactive Intelligent Systems (TIIS), 9(2–3): 12:1–12:??, April 2019. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3213769.

Sovrano:2022:GUC

[SV22] Francesco Sovrano and Fabio Vitali. Generating usercentred explanations via illocutionary question answering: From philosophy to interfaces. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(4): 26:1–26:??, December 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3519265.

Sappelli:2016:AIA

[SVK16] Maya Sappelli, Suzan Verberne, and Wessel Kraaij. Adapting the interactive activation model for context recognition and identification. ACM Transactions on Interactive Intelligent Systems (TIIS), 6(3):22:1–22:??, October 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Springer:2020:PDW

[SW20] Aaron Springer and Steve Whittaker. Progressive disclosure: When, why, and how do users want algorithmic transparency information? ACM Transactions on Interactive Intelligent Systems (TIIS), 10(4): 29:1–29:32, December 2020. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3374218.

Starke:2021:PEE

[SWS21] Alain Starke, Martijn Willemsen, and Chris Snijders. Promoting energy-efficient behavior by depicting social norms in a recommender interface. ACM Transactions on Interactive Intelligent Systems (TIIS), 11 (3–4):30:1–30:32, December 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3460005.

Sartori:2015:AAP

[SYS^{+}15] Andreza Sartori, Victoria Yanulevskaya, Almila Akdag Salah, Jasper Uijlings, Elia Bruni, and Nicu Sebe. Affective analysis of professional and amateur abstract paintings using statistical analysis and art theory. ACM Transactions on Interactive Intelligent Systems (TIIS), 5(2): 8:1–8:??, July 2015. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Tsai:2020:ESR

[TB20] Chun-Hua Tsai and Peter Brusilovsky. Exploring social recommendations with visual diversity-promoting interfaces. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(1): 5:1–5:34, January 2020. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3231465.

Taib:2020:PSD

[TBK^{+}20] Ronnie Taib, Shlomo Berkovsky, Irena Koprinska, Eileen Wang, Yucheng Zeng, and Jingjie Li. Personality sensing: Detection of personality traits using physiological responses to image and video stimuli. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(3): 18:1–18:32, November 2020. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3357459.

Thebault:2013:ESP

[TDBL13] Pierrick Thebault, Dominique Decotter, Mathieu Boussard, and Monique Lu. Embodying services into physical places: Toward the design of a mobile environment browser. ACM Transactions on Interactive Intelligent Systems (TIIS), 3 (2):8:1–8:??, July 2013. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Tran:2021:HRS

[TFT21] Thi Ngoc Trang Tran, Alexander Felfernig, and Nava Tintarev. Humanized recommender systems: Stateof-the-art and research issues. ACM Transactions on Interactive Intelligent Systems (TIIS), 11(2):9:1–9:41, July 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3446906.

Todi:2020:IGL

[TJLO20] Kashyap Todi, Jussi Jokinen, Kris Luyten, and Antti Oulasvirta. Individualising graphical layouts with predictive visual search models. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(1):9:1–9:24, January 2020. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL

https://dl.acm.org/doi/
10.1145/3241381.

Tintarev:2016:ISI

[TOF16] Nava Tintarev, John O’donovan, and Alexander Felfernig. Introduction to the special issue on human interaction with artificial advice givers. ACM Transactions on Interactive Intelligent Systems (TIIS), 6(4):26:1–26:??, December 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Thomason:2019:CAV

[TRO^{+}19] John Thomason, Photchara Ratsamee, Jason Orlosky, Kiyoshi Kiyokawa, Tomohiro Mashita, Yuki Uranishi, and Haruo Takemura. A comparison of adaptive view techniques for exploratory 3D drone teleoperation. ACM Transactions on Interactive Intelligent Systems (TIIS), 9 (2–3):17:1–17:??, April 2019. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3232232.

Tanaka:2016:TSC

[TSN^{+}16] Hiroki Tanaka, Sakti Sakriani, Graham Neubig, Tomoki Toda, Hideki Negoro, Hidemi Iwasaka, and Satoshi Nakamura. Teaching social communication skills through human-agent interaction. ACM Transactions on Interactive Intelligent Systems (TIIS), 6(2):18:1–18:??, August 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Taranta:2015:EBC

[TSSL15] Eugene M. Taranta II, Thaddeus K. Simons, Rahul Sukthankar, and Joseph J. Laviola, Jr. Exploring the benefits of context in 3D gesture recognition for gamebased virtual environments. ACM Transactions on Interactive Intelligent Systems (TIIS), 5(1):1:1–1:??, March 2015. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Tang:2022:SOA

[TTL^{+}22] Tan Tang, Junxiu Tang, Jiewen Lai, Lu Ying, Yingcai Wu, Lingyun Yu, and Peiran Ren. SmartShots: an optimization approach for generating videos with data visualizations embedded. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(1):4:1–4:21, March 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3484506.

Turkay:2021:SII

[TVA^{+}21] Cagatay Turkay, Tatiana Von Landesberger, Daniel Archambault, Shixia Liu, and Remco Chang. Special issue on interactive visual analytics for making explainable and accountable decisions. ACM Transactions on Interactive Intelligent Systems (TIIS), 11(3–4):17:1–17:4, December 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3471903.

Taranta:2016:DPB

[TVCL16] Eugene M. Taranta II, Andrés N. Vargas, Spencer P. Compton, and Joseph J. Laviola, Jr. A dynamic penbased interface for writing and editing complex mathematical expressions with math boxes. ACM Transactions on Interactive Intelligent Systems (TIIS), 6 (2):13:1–13:??, August 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Vanderdonckt:2020:EDS

[VBCC20] Jean Vanderdonckt, Sara Bouzit, Gaëlle Calvary, and Denis Chêne. Exploring a design space of graphical adaptive menus: Normal vs. small screens. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(1): 2:1–2:40, January 2020. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3237190.

vandeGarde-Perik:2013:AIO

[vdGPOvB^{+}13] Evelien van de Garde-Perik, Serge Offermans, Koen van Boerdonk, Kars-Michiel Lenssen, and Elise van den Hoven. An analysis of input-output relations in interaction with smart tangible objects. ACM Transactions on Interactive Intelligent Systems (TIIS), 3(2): 9:1–9:??, July 2013. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Vezzani:2015:GPS

[VLP^{+}15] Roberto Vezzani, Martino Lombardi, Augusto Pieracci, Paolo Santinelli, and Rita Cucchiara. A generalpurpose sensing floor architecture for human-environment interaction. ACM Transactions on Interactive Intelligent Systems (TIIS), 5(2): 10:1–10:??, July 2015. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

VanBerkel:2022:IRF

[VOA^{+}22] Niels Van Berkel, Jeremy Opie, Omer F. Ahmad, Laurence Lovat, Danail Stoyanov, and Ann Blandford. Initial responses to false positives in AI-supported continuous interactions: a colonoscopy case study. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(1):2:1–2:18, March 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3480247.

Vainio-Pekka:2023:REA

[VPAJ^{+}23] Heidi Vainio-Pekka, Mamia Ori-Otse Agbese, Marianna Jantunen, Ville Vakkuri, Tommi Mikkonen, Rebekah Rousi, and Pekka Abrahamsson. The role of explainable AI in the research field of AI ethics. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(4): 26:1–26:??, December 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3599974.

Verbert:2016:AVU

[VPB16] Katrien Verbert, Denis Parra, and Peter Brusilovsky. Agents vs. users: Visual recommendation of research talks with multiple dimension of relevance. ACM Transactions on Interactive Intelligent Systems (TIIS), 6 (2):11:1–11:??, August 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Vig:2012:TGE

[VSR12] Jesse Vig, Shilad Sen, and John Riedl. The tag genome: Encoding community knowledge to support novel interaction. ACM Transactions on Interactive Intelligent Systems (TIIS), 2(3): 13:1–13:??, September 2012. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Wetzler:2013:CPM

[WBL^{+}13] Philipp Wetzler, Steven Bethard, Heather Leary, Kirsten Butcher, Soheil Danesh Bahreini, Jin Zhao, James H. Martin, and Tamara Sumner. Characterizing and predicting the multifaceted nature of quality in educational Web resources. ACM Transactions on Interactive Intelligent Systems (TIIS), 3 (3):15:1–15:??, October 2013. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Wang:2020:MDS

[WCS20] Ruijie Wang, Liming Chen, and Ivar Solheim. Modeling dyslexic students’ motivation for enhanced learning in e-learning systems. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(3):21:1–21:34, November 2020. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3341197.

Wenskovitch:2024:TAA

[WDN24] John Wenskovitch, Michelle Dowling, and Chris North. Toward addressing ambiguous interactions and inferring user intent with dimension reduction and clustering combinations in visual analytics. ACM Transactions on Interactive Intelligent Systems (TIIS), 14(1): 2:1–2:??, March 2024. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3588565.

Wang:2016:ART

[WES16] Weiyi Wang, Valentin Enescu, and Hichem Sahli. Adaptive real-time emotion recognition from body movements. ACM Transactions on Interactive Intelligent Systems (TIIS), 5 (4):18:1–18:??, January 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Wang:2018:TBM

[WHLZ18] Yue Wang, Laura R. Humphrey, Zhanrui Liao, and Huanfei Zheng. Trust-based multi-robot symbolic motion planning with a human-in-the-loop. ACM Transactions on Interactive Intelligent Systems (TIIS), 8(4): 31:1–31:??, November 2018. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3213013.

Wagner:2018:MHR

[WRH18] Alan R. Wagner, Paul Robinette, and Ayanna Howard. Modeling the human-robot trust phenomenon: a conceptual framework based on risk. ACM Transactions on Interactive Intelligent Systems (TIIS), 8(4): 26:1–26:??, November 2018. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3152890.

Wang:2018:VRI

[WSL^{+}18] Yong Wang, Conglei Shi, Liangyue Li, Hanghang Tong, and Huamin Qu. Visualizing research impact through citation data. ACM Transactions on Interactive Intelligent Systems (TIIS), 8(1):5:1–5:??, March 2018. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Wang:2023:WBH

[WWB^{+}23] Clarice Wang, Kathryn Wang, Andrew Y. Bian, Rashidul Islam, Kamrun Naher Keya, James Foulds, and Shimei Pan. When biased humans meet debiased AI: a case study in college major recommendation. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(3):17:1–17:??, September 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3611313.

Wang:2022:EEA

[WY22] Xinru Wang and Ming Yin.

Effects of explanations in AI-assisted decision making: Principles and comparisons. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(4): 27:1–27:??, December 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3519266.

Wade:2016:GCA

[WZB^{+}16] Joshua Wade, Lian Zhang, Dayi Bian, Jing Fan, Amy Swanson, Amy Weitlauf, Medha Sarkar, Zachary Warren, and Nilanjan Sarkar. A gaze-contingent adaptive virtual reality driving environment for intervention in individuals with autism spectrum disorders. ACM Transactions on Interactive Intelligent Systems (TIIS), 6(1):

3:1–3:??, May 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Xu:2016:SYS

[XZY16] Tian (Linger) Xu, Hui Zhang, and Chen Yu. See you see me: The role of eye contact in multimodal human-robot interaction. ACM Transactions on Interactive Intelligent Systems (TIIS), 6(1): 2:1–2:??, May 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Yang:2021:TRA

[Yan21] Qiang Yang. Toward responsible AI: an overview of federated learning for usercentered privacy-preserving computing. ACM Transactions on Interactive Intelligent Systems (TIIS), 11 (3–4):32:1–32:22, December 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3485875.

Yan:2023:IPT

[YC23] Dongning Yan and Li Chen. The influence of personality traits on user interaction with recommendation interfaces. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(1): 3:1–3:??, March 2023. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3558772.

Yan:2017:EAR

[YDL^{+}17] Shuo Yan, Gangyi Ding, Hongsong Li, Ningxiao Sun, Zheng Guan, Yufeng Wu, Longfei Zhang, and Tianyu Huang. Exploring audience response in performing arts with a brain-adaptive digital performance system. ACM Transactions on Interactive Intelligent Systems (TIIS), 7(4):16:1–16:??, December 2017. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Young:2014:DET

[YIS^{+}14] James E. Young, Takeo Igarashi, Ehud Sharlin, Daisuke Sakamoto, and Jeffrey Allen. Design and evaluation techniques for authoring interactive and stylistic behaviors. ACM Transactions on Interactive Intelligent Systems (TIIS), 3(4): 23:1–23:??, January 2014. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Yordanova:2016:PSD

[YK16] Kristina Yordanova and Thomas Kirste. A process for systematic development of symbolic models for activity recognition. ACM Transactions on Interactive Intelligent Systems (TIIS), 5 (4):20:1–20:??, January 2016.

CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Yalcin:2023:IIP

[YLC23] Özge Nilay Yalçın, Sébastien Lallé, and Cristina Conati. The impact of intelligent pedagogical agents’ interventions on student behavior and performance in openended game design environments. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(3): 11:1–11:??, September 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3578523.

Yazdani:2012:ARB

[YLVE12] Ashkan Yazdani, Jong-Seok Lee, Jean-Marc Vesin, and Touradj Ebrahimi. Affect recognition based on physiological changes during the watching of music videos. ACM Transactions on Interactive Intelligent Systems (TIIS), 2(1):7:1–7:??, March 2012. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Yu:2013:ISI

[YN13] Lixiu Yu and Jeffrey V. Nickerson. An Internet-scale idea generation system. ACM Transactions on Interactive Intelligent Systems (TIIS), 3 (1):2:1–2:??, April 2013. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Yagi:2022:GFR

[YNK^{+}22] Takuma Yagi, Takumi Nishiyasu, Kunimasa Kawasaki, Moe Matsuki, and Yoichi Sato. GO-Finder: a registrationfree wearable system for assisting users in finding lost hand-held objects. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(4):35:1–35:??, December 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3519268.

Yang:2016:SUS

[YPL^{+}16] Yi Yang, Shimei Pan, Jie Lu, Mercan Topkara, and Yangqiu Song. The stability and usability of statistical topic models. ACM Transactions on Interactive Intelligent Systems (TIIS), 6 (2):14:1–14:??, August 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Ye:2015:UUS

[YSD15] Juan Ye, Graeme Stevenson, and Simon Dobson. US-MART: an unsupervised semantic mining activity recognition technique. ACM Transactions on Interactive Intelligent Systems (TIIS), 4 (4):16:1–16:??, January 2015.

CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Yang:2015:QSM

[YT15] Yi-Hsuan Yang and Yuan-Ching Teng. Quantitative study of music listening behavior in a Smartphone context. ACM Transactions on Interactive Intelligent Systems (TIIS), 5(3):14:1–14:??, October 2015. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Yousefi:2024:EFM

[YVA^{+}24] Zeinab R. Yousefi, Tung Vuong, Marie AlGhossein, Tuukka Ruotsalo, Giulio Jaccuci, and Samuel Kaski. Entity footprinting: Modeling contextual user states via digital activity monitoring. ACM Transactions on Interactive Intelligent Systems (TIIS), 14(2):9:1–9:??, June 2024. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3643893.

Yamazaki:2016:ITN

[YYI^{+}16] Keiichi Yamazaki, Akiko Yamazaki, Keiko Ikeda, Chen Liu, Mihoko Fukushima, Yoshinori Kobayashi, and Yoshinori Kuno. “I’ll Be There Next”: a multiplex care robot system that conveys service order using gaze gestures. ACM Transactions on Interactive Intelligent Systems (TIIS), 5(4): 21:1–21:??, January 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Zhou:2023:TBP

[ZB23] Michelle Zhou and Shlomo Berkovsky. 2022 TiiS best paper announcement. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(3):19:1–19:??, September 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3615590.

Zamborlin:2014:FGI

[ZBGD14] Bruno Zamborlin, Frederic Bevilacqua, Marco Gillies, and Mark D’inverno. Fluid gesture interaction design: Applications of continuous recognition for the design of modern gestural interfaces. ACM Transactions on Interactive Intelligent Systems (TIIS), 3(4):22:1–22:??, January 2014. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Zhang:2018:ERC

[ZCC^{+}18] Amy X. Zhang, Jilin Chen, Wei Chai, Jinjun Xu, Lichan Hong, and Ed Chi. Evaluation and refinement of clustered search results with the crowd. ACM Transactions on Interactive Intelligent Systems (TIIS), 8(2):

14:1–14:??, July 2018. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Zhang:2021:MMI

[ZCC21] Yu Zhang, Bob Coecke, and Min Chen. MI3: Machineinitiated intelligent interaction for interactive classification and data reconstruction. ACM Transactions on Interactive Intelligent Systems (TIIS), 11(3–4):18:1–18:34, December 2021. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3412848.

Zanzotto:2017:YLT

[ZF17] Fabio Massimo Zanzotto and Lorenzo Ferrone. Have you lost the thread? Discovering ongoing conversations in scattered dialog blocks. ACM Transactions on Interactive Intelligent Systems (TIIS), 7 (2):9:1–9:??, July 2017. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Zhang:2024:KLB

[ZFM^{+}24] Rui Zhang, Christopher Flathmann, Geoff Musick, Beau Schelble, Nathan J. McNeese, Bart Knijnenburg, and Wen Duan. I know this looks bad, but I can explain: Understanding when AI should explain actions in Human–AI teams. ACM Transactions on Interactive Intelligent Systems (TIIS), 14(1):6:1–6:??, March 2024. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3635474.

Zhang:2016:BTE

[ZGZ^{+}16] Cheng Zhang, Anhong Guo, Dingtian Zhang, Yang Li, Caleb Southern, Rosa I. Arriaga, and Gregory D. Abowd. Beyond the touchscreen: an exploration of extending interactions on commodity Smartphones. ACM Transactions on Interactive Intelligent Systems (TIIS), 6 (2):16:1–16:??, August 2016. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Zhou:2020:ITS

[Zho20] Michele X. Zhou. Introduction to the TiiS special column. ACM Transactions on Interactive Intelligent Systems (TIIS), 10(4): 25:1, December 2020. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3427592.

Zhou:2021:ISC

[Zho21] Michelle X. Zhou. Introduction to the special column for human-centered artificial intelligence. ACM Transactions on Interactive Intelligent Systems (TIIS), 11(3– 4):31:1, December 2021. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3490553.

Zhou:2022:EIT

[Zho22] Michelle X. Zhou. Editorial introduction to TiiS special category article: Practitioners’ toolbox. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(1): 7:1, March 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3519381.

Zini:2022:ACT

[ZLG22] Floriano Zini, Fabio Le Piane, and Mauro Gaspari. Adaptive cognitive training with reinforcement learning. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(1):3:1– 3:29, March 2022. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3476777.

Zhang:2017:EEI

[ZLW17] Ting Zhang, Yu-Ting Li, and Juan P. Wachs. The effect of embodied interaction in visual-spatial navigation. ACM Transactions on Interactive Intelligent Systems (TIIS), 7(1):3:1–3:??, March 2017. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic).

Zhou:2019:TVA

[ZMLY19] Michelle X. Zhou, Gloria Mark, Jingyi Li, and Huahai Yang. Trusting virtual agents: The effect of personality. ACM Transactions on Interactive Intelligent Systems (TIIS), 9(2–3): 10:1–10:??, April 2019. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3232077.

Zorrilla:2022:MNC

[ZT22] Asier López Zorrilla and M. Inés Torres. A multilingual neural coaching model with enhanced long-term dialogue structure. ACM Transactions on Interactive Intelligent Systems (TIIS), 12(2): 16:1–16:47, June 2022. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

dl.acm.org/doi/10.1145/
3487066.

Zhang:2024:SBO

[ZTD^{+}24] Yu Zhang, Martijn Tennekes, Tim De Jong, Lyana Curier, Bob Coecke, and Min Chen. Simulation-based optimization of user interfaces for quality-assuring machine learning model predictions. ACM Transactions on Interactive Intelligent Systems (TIIS), 14(1):

1:1–1:??, March 2024. CO-DEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3594552.

Zoller:2023:XVA

[ZTSH23] Marc-André Zöller, Waldemar Titov, Thomas Schlegel, and Marco F. Huber. XAutoML: a visual analytics tool for understanding and validating automated machine learning. ACM Transactions on Interactive Intelligent Systems (TIIS), 13(4): 28:1–28:??, December 2023. CODEN ???? ISSN 2160-6455 (print), 2160-6463 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3625240.